### v1

# M2D2: Exploring General-purpose Audio-Language Representations Beyond CLAP

Abstract
Contrastive language-audio pre-training (CLAP) has addressed audio-language tasks such as audio-text retrieval by aligning audio and text in a common feature space. While CLAP addresses general audio-language tasks, its audio features do not generalize well in audio tasks. In contrast, self-supervised learning (SSL) models learn general-purpose audio features that perform well in diverse audio tasks. We pursue representation learning that can be widely used in audio applications and hypothesize that a method that learns both general audio features and CLAP features should achieve our goal, which we call a general-purpose audio-language representation. To implement our hypothesis, we propose M2D2, a second-generation masked modeling duo (M2D) that combines an SSL M2D and CLAP. M2D2 learns two types of features using two modalities (audio and text) in a two-stage training process. It also utilizes advanced LLM-based sentence embeddings in CLAP training for powerful semantic supervision. In the first stage, M2D2 learns generalizable audio features from M2D and CLAP, where CLAP aligns the features with the fine LLM-based semantic embeddings. In the second stage, it learns CLAP features using the audio features learned from the LLM-based embeddings. Through these pre-training stages, M2D2 should enhance generalizability and performance in its audio and CLAP features. Experiments validated that M2D2 achieves effective general-purpose audio-language representation, highlighted with SOTA fine-tuning mAP of 49.0 for AudioSet, SOTA performance in music tasks, and top-level performance in audio-language tasks.
Index Terms:
Audio representation learning, general-purpose audio-language representation, masked modeling duo, CLAP.I Introduction
Contrastive language-image pre-training (CLIP)[1] has enabled the alignment of two modalities and has significantly impacted diverse domains. In the audio domain, contrastive language-audio pre-training (CLAP) has extended CLIP to audio[2][3][4][5], followed by various advancements[6][7][8][9][10][11] that have exhibited superior performance for applications such as zero-shot classification and audio captioning[7][12][13]. CLAP can map audio and captions into features aligned in common feature space; therefore, these features can be effective for various downstream application tasks since many task settings are typically related to language, such as inference of class labels.
While the CLAP models have exhibited versatility in audio-language tasks, their audio features do not generalize well in conventional audio tasks because they typically use audio encoders based on supervised learning models. The state-of-the-art (SOTA) audio representations are self-supervised learning (SSL) models known to generalize to diverse downstream tasks[14]. In contrast, the supervised learning models, and thus the CLAP models, suffer on tasks different from the pre-training, as we reported in [11][15]. For example, they perform well in ESC-50[16], an environmental sound classification similar to the pre-trained AudioSet[17] but poorly in VoxCeleb1[18], a speaker identification task. However, some studies have applied CLAP audio features to a task where SSL audio features may perform well, such as [19][20].
We pursue representation learning that can be widely effective in diverse audio applications and hypothesize that a generally useful method would be one that learns both general-purpose audio features and CLAP features; we call it a general-purpose audio-language representation. Since the current mainstream audio representation learning methods are based on SSL with a masked prediction objective, we think combining CLAP and masked prediction-based SSL should achieve our goal. However, as contrastive learning in CLAP is memory-intensive, naively combining it with SSL requires large computational resources.
On the other hand, to learn the best of CLAP, a SOTA sentence embedding model based on a large-language model (LLM) would be a good candidate as a CLAP text encoder. In the natural language processing (NLP) domain, sentence embedding models have recently progressed to exploit an LLM[21]. LLMs have demonstrated overwhelming ability in language-related tasks[22] and have the potential to map audio captions into better semantic features based on knowledge accumulated in huge parameters learned from massive data.
To implement a general-purpose audio-language representation and achieve its highest potential, we propose M2D2, which extends a SOTA masked prediction-based SSL, M2D[23], combines it with a CLAP objective, and exploits the powerful LLM-based sentence embeddings, as shown in Fig. 1. It uses a two-stage pre-training strategy that suppresses required computational resources, thus enabling multitasking of M2D and CLAP, contrastive learning with larger batch size, and even learning from LLM-based sentence embeddings.
The two-stage pre-training strategy enables learning the best features while avoiding substantial computational costs. In the first stage, M2D2 learns from a multitask of M2D and CLAP and uses LLM-based text features for strong semantic supervision, aiming to learn the best general-purpose audio features that can also be projected to an effective CLAP feature. In the second stage, M2D2 pre-trains a non-LLM text encoder to learn effective CLAP text features that align better with the learned audio features. We aim to exploit the LLM-based knowledge in the sentence embeddings through the stages while avoiding the cost of huge computation by training either audio or text encoder in a stage.
The experiments validate the effectiveness of M2D2 in achieving our goal of a general-purpose audio-language representation. The learned audio feature exhibits SOTA performance in fine-tuning, and the CLAP features demonstrate top performance in audio-language tasks. For a successful pre-training by multitasking M2D and CLAP, we find it essential to balance the loss between M2D and CLAP to control the powerful LLM-based semantic supervision. The audio projector capacity, LLM-based text encoder, and two-stage training strategy designs are also crucial, which we demonstrate in the ablation studies in Section IV-D.
Our contributions are:
A general-purpose audio-language representation that provides CLAP features for tasks where audio-language alignment is beneficial and general-purpose audio features for other conventional audio tasks.
A new pre-training method, M2D2, that combines M2D and CLAP using LLM-based sentence embeddings and provides SOTA general-purpose audio features and top-performing CLAP features, achieving a general-purpose audio-language representation.
Empirical findings on combining SSL and CLAP in ablation studies and publicly available code with pre-trained weights111https://github.com/nttcslab/m2d for future studies.
The first study to show that an LLM-based sentence embedding model from the NLP domain helps learn effective audio and CLAP features.
II Related Work
II-A Relationship with our previous work
This study redefines and extends our previous work, M2D-CLAP[11]. M2D2 differs from M2D-CLAP in improved audio and CLAP features achieved by the newly employed two-stage pre-training strategy, improved audio projector based on a transformer encoder, and the use of LLM-based sentence embeddings (Section III-B). To explore its effectiveness further, we also conducted experiments on audio-text retrieval (Section IV-C2), audio captioning (Section IV-C3), and ablation studies (Section IV-D) in addition to audio transfer learning (Section IV-B) and zero-shot classification (Section IV-C1).
To make the difference from the final M2D2 clear, we specifically denote M2D-CLAP as M2D20- in the experiments because it is a variant of M2D2-, the first-stage pre-trained M2D2.
II-B Conventional audio representation models
Since the release of the large-scale dataset AudioSet[17], various supervised learning models such as PANNs[24], AST[25], and HTS-AT[26] that utilize class labels have followed. In contrast, SSL methods such as COLA[27] and BYOL-A[15] are general-purpose audio signal representations that do not depend on task supervision, such as labels. Masked prediction-based SSLs have progressed and demonstrated remarkable general-purpose performance. SSAST[28], MAE-AST[29], MSM-MAE[30], and Audio-MAE[31] learn by reconstructing the masked input, while M2D and ATST[32] learn by predicting the masked features. More advanced models include BEATs[33], which learn by predicting distilled and tokenized labels, and CED[34], which distills multiple large models. While they perform well on conventional audio tasks, their features do not align with languages.
II-C CLAP models
Inspired by CLIP[1], the early CLAP models AudioCLIP[2] and Wav2CLIP[3] learn audio features that align with the pre-trained CLIP multimodal feature. CLAP[5][6] and LAION-CLAP[4] adapt CLIP to audio, WavCaps[7] improves performance with their dataset, and MGA-CLAP[35] further improves fine-grained audio-language alignment. While they can provide an audio feature as a part of CLAP features, they do not generalize well on conventional audio tasks, as reported in [11][15].
Methods that take a similar approach to M2D2 are SLIP[36] in the image domain (combines SSL and CLIP), FLAP[8] (combines MAE[37] and CLAP), SupMAM-CLAP[38] (combines Audio-MAE and distillation of a CLAP model), Cacophony[10] (combines Audio-MAE, CLAP, and a captioning model in a two-stage training), and CoLLAT[9] (CLAP with frozen text encoders). Unlike these approaches, we combine M2D, CLAP, and the recent LLM-based sentence embeddings in a two-stage pre-training to learn a general-purpose audio-language representation that provides generalizable audio and CLAP features.
III Proposed Method
We propose M2D2, which extends a masked prediction-based SSL method, M2D, to combine it with CLAP to learn general-purpose audio-language representations effective for a broad range of conventional audio and audio-language tasks.
III-A Background: Masked Modeling Duo
M2D is an SSL framework applicable to 2D structured data input such as audio spectrograms, and it pre-trains Vision Transformer[39] (ViT) with a masked prediction objective to learn a general-purpose representation.
III-A1 Pre-training
As shown in Fig. 2(a), M2D consists of two networks, the online and the target, and learns representations by predicting the target output features using the online output features. It takes a spectrogram with a predefined size (e.g., 80 frequency bins and 608 time steps) as the input, which is split into patches (e.g., ), mapped by a learnable linear layer (patch embedding layer), and handled as a series (e.g., patches). M2D then adds positional encoding to patches and randomly selects a number of patches according to a masking ratio as masked patches (e.g., 70% of the input patches) and the rest as visible patches (e.g., the remaining 30%).
The online network with a set of weights encodes using the online audio encoder into the audio features . It concatenates the learnable mask tokens to , adds the position encoding , and inputs them to the predictor to predict the target features as . It then outputs the prediction result of the masked patch features, where is the set of masked patch indices.
The target network, defined by parameter , outputs the features and standardizes them to the final target output . The loss is calculated using the online prediction against the actual target output as a training signal by the mean square error (MSE) of -normalized and :
where denotes the inner product.
The M2D framework updates only to minimize the loss , as depicted by the stop-gradient in Fig. 2(a), and updates as an exponential moving average of with a decay rate . M2D exploits the momentum encoder to learn effective representations from the target network.
III-A2 Transfer learning
After the pre-training, only the audio encoder is transferred from the online network in downstream tasks. As the encoder accepts the input no longer than the predefined duration (e.g., 608 time steps) in the pre-training stage, the inference procedure handles variable-length audio inputs in two ways: one is by cutting the input with the predefined duration and repeatedly encoding the chunks; the other is by interpolating the positional encoding to the input audio duration (i.e., stretching the acceptable input duration). The latter is used for fine-tuning the encoder to a task with an input duration similar to the pre-training and the former is used in all other cases, such as linear evaluation.
According to the task requirements, the encoder outputs a feature for each patch , which is summarized into a time frame-level feature or an audio clip-level feature . The following pseudo-code calculates for each time frame:
where , is batch size, is the number of patches along frequency, is the number of patches along time, is a feature dimension, and is the calculation result. The audio clip-level feature is calculated by averaging over time: . For example, becomes a 3840-d feature, where and .
III-A3 M2D for X
In addition to SSL, M2D provides a framework for learning a specialized representation in application X (M2D-X) by extending M2D to add an application-specific task and form a multitask of M2D and the added task in parallel. During the pre-training, the intermediate outputs and in M2D are concatenated and fed into the additional task, and the loss of M2D and the additional task are combined to train the entire network. As a result, the learned representation reflects both objectives from M2D and the added task.
III-B Proposed M2D2: M2D Second Generation
M2D2 is a second-generation M2D, an M2D-X variant that adds a CLAP training objective to implement a general-purpose audio-language representation. As shown in Fig. 2, it takes a multitask approach comprising (a) M2D and (b) CLAP that learns from paired audio and caption data. The CLAP extension has a semantic network branch that conducts contrastive learning among audio and text features to align them in a semantic feature space.
M2D2 employs a two-stage training strategy to learn the best of audio and CLAP features under a multitask of M2D and CLAP without the cost of huge computational resources. In the first stage, it learns effective audio features from M2D that also jointly learns from powerful LLM-based sentence embeddings by aligning the audio features with them through CLAP, aiming to achieve the best audio features that are also effective as CLAP audio features. In the second stage, it learns effective CLAP text features that align with the learned audio features in a semantic feature space. Through these stages, M2D2 learns effective audio and CLAP features.
III-B1 First stage pre-training
In the first stage, the captions are encoded into embeddings in advance using the LLM-based text encoder, and the training process takes audio for M2D input while taking text embeddings for the corresponding caption. M2D then conducts its training process and feeds the audio encoder’s output feature to the semantic network. While the original M2D-X framework provides in addition to , M2D2 uses only for training stability. The audio projector converts the audio features into semantic features . The text embeddings are used as they are or mapped to the same dimension as using a learnable linear layer and handled as text semantic features . The training process does not involve the LLM-based sentence embedding model, which requires a large GPU memory.
This stage aims to train an effective audio encoder, so it is crucial that the training signal, semantic feature , precisely represents the semantics of the text. We use a powerful LLM-based sentence embedding model and distill its semantic knowledge into the audio encoder through CLAP. The audio projector is a transformer encoder to summarize so that the can better represent the semantics of an audio input.
The training process calculates the CLAP loss using the cosine similarity between and :
where is the semantic feature of the th audio batch sample, and is the semantic feature of the th caption batch sample. The CLAP loss is the average of the NT-Xent[40] losses calculated along the audio and caption axes:
where is the number of batch samples, and is a learnable temperature parameter. We follow CLIP [1] to initialize with 0.07 and clip it to prevent scaling the logits by more than 100 for training stability.
The entire loss combines and :
where the loss weights and control the contribution.
III-B2 Second stage pre-training
The second stage pre-trains the text encoder and audio projector using CLAP alone and freezes the audio encoder weights so that we preserve effective audio features learned in the first stage, as shown in Fig. 3. To facilitate the CLAP training using a precise loss, we make the audio feature for an audio input consistent by using a low masking ratio and inputting the entire audio sample as is.
In this stage, we employ a modest capacity model, such as BERT, as the text encoder to maintain effective contrastive learning with a sufficient batch size. The training process masks the audio patches with a low masking ratio (e.g., 30%), the frozen audio encoder encodes input patches into the features , and the audio projector summarizes them into the semantic features . The text encoder encodes captions into the semantic features , and the clap loss is calculated between and , which is the loss for the second stage.
III-B3 M2D2 transfer learning
Once pre-trained, we transfer the audio encoder, audio projector, and text encoder as models constituting a general-purpose audio-language representation to the downstream tasks. The audio encoder outputs audio features of the input audio, the audio projector converts into CLAP audio features , and the text encoder encodes captions into CLAP text features . While the audio feature is useful for conventional audio tasks, and work as CLAP features in audio-language tasks such as zero-shot classification and audio-text retrieval.
III-B4 Additional pre-training stages
While the two training stages enable the learning of general-purpose features, the experiments show that these features still leave a gap with top performance in existing benchmarks if they are not fine-tuned. Therefore, to facilitate the comparison with previous studies, we also use two additional stages to bring the experimental conditions closer to those of existing studies.
Stage 1.1: AudioSet fine-tuning
This stage fine-tunes the first-stage pre-trained audio encoder to AudioSet under a supervised setting using labels. This additional process corresponds to adopting the AudioSet fine-tuned audio encoders in the conventional CLAP models[4][5][6]. Experiments in Sections IV-C2 and IV-C3 show that this fine-tuning step is advantageous with AudioCaps[41] (a subset of AudioSet).
Stage 2.1: WavCaps pre-training
This stage further pre-trains the second stage audio projector and text encoder to refine the CLAP features using the same dataset as in WavCaps[7] and with no masking of the audio input to learn the best text/audio alignment. We show that this process improves performance in audio-language tasks in Section IV-C.
III-B5 Training design and efficiency
Though other training strategies are possible, we pre-train the audio encoder first because we aim to exploit the knowledge from the LLM for both audio and text features. The M2D2 first stage enables that, unlike typical CLAP models that use a pre-trained audio encoder solely learned from audio data. To practically train a CLAP, we need pre-trained audio/text encoders, which is why we train the text encoder in the second stage. To enable effective training without a huge computing budget, we train encoders separately and use the LLM offline.
We utilize an audio projector only, unlike previous CLAP, which typically uses both audio and text projectors. An audio projector is crucial to bridge the gap between the descriptive general audio and abstract (semantic) CLAP features so that they can be learned in a single model. A text projector degrades performance in M2D2, as verified in Section IV-D3.
The staged pre-training allows for training under fewer computing resources, in addition to learning a well-performing representation. M2D2 learns from LLM features in the first stage, while the features are constant for the same captions. Therefore, M2D2 can substantially reduce resource requirements by pre-computing and retaining features for all captions and referring to pre-computed features instead of loading an LLM instance during training. In contrastive learning, a large batch size is critical for sufficient training. While FLAP, a similar method, requires GPUs to achieve a batch size of 2304, M2D2 can conduct training with 4 A100 GPUs (80GB) for a batch size of 2048.
IV Experiments
To assess the effectiveness of M2D2 representations, we evaluated the audio features on conventional audio tasks under settings of frozen audio features (Sections IV-B1 and IV-B2) and fine-tuning audio features (Section IV-B3) and the CLAP features using three tasks: zero-shot classification (Section IV-C1), audio-text retrieval (Section IV-C2), and audio captioning (Section IV-C3).
We also studied how the pre-training design and parameters impact the final performance (Section IV-D), compared the CLAP features using visualizations (Section IV-D4), and show how M2D2 focuses on the audio patches to summarize CLAP audio features (Section IV-D4). In the following sections, we refer to the M2D2 models as follows:
M2D2-: First stage pre-trained representation. Combination of pre-trained audio encoder/audio projector and off-the-shelf LLM-based sentence embedding model as a text encoder.
M2D2: Second stage pre-trained representation. Combination of the first-stage audio encoder and the second-stage pre-trained audio projector/text encoder.
M2D2+: Second stage plus stage 2.1 pre-trained representation. Combination of the first-stage audio encoder and the stage 2.1 pre-trained audio projector/text encoder.
M2D2: Fine-tuned representation to the benchmarks. Combination of the audio encoder fine-tuned in stage 1.1 and the stage 2.1 pre-trained audio projector/text encoder.
M2D20-: The M2D-CLAP[11] model as an ablation variant of M2D2- for using a non-LLM text encoder and AudioSet only for pre-training audio data.
IV-A Experimental setup
IV-A1 Basic setup
We followed M2D for the basic setup and implemented M2D2 on top of the M2D codebase as a variant of M2D-X[23]. Specifically, the audio encoder is ViT-base[39] with a 768-d feature, input audio duration of 6 s, and patch size of . The audio projector is a single-block transformer encoder with a 768-d feature, a single head, a 768-d feed-forward network latent feature, and a class token. It inputs audio features encoded by the audio encoder and outputs the class token feature as a summarized CLAP audio feature.
The audio samples were preprocessed to a log-scaled mel spectrogram with a sampling frequency of 16,000 Hz, window size of 25 ms, hop size of 10 ms, and 80 mel-spaced frequency bins in the range of 50 to 8000 Hz. We standardized spectrograms with the statistics of the first stage pre-training dataset (e.g., average of and standard deviation of ).
IV-A2 Pre-training setup
The first stage
In the first stage, we used the M2D masking ratio of 0.7, and all other settings were the same as in M2D, such as 300 epochs, 20 warm-up epochs, batch size of 2048, the base learning rate of , and the linearly interpolated EMA decay rate starting from 0.99995 to 0.99999 at the end. The loss weights were and .
We used a generalist text embedding model NV-Embed-v2[21][42] with 4096-d features as a text encoder from Hugging Face222https://huggingface.co/nvidia/NV-Embed-v2, built on top of an LLM, Mistral 7B[22]. NV-Embed-v2 exhibits SOTA performance on the Massive Text Embedding Benchmark (MTEB)[43]. We used a learnable linear layer to map the 4096-d output feature to 768 dimensions.
The second stage
In the second stage, we used the M2D masking ratio of 0.3, 30 epochs, five warm-up epochs, batch size of 2048, and the base learning rate of . We used BERT-base[44] as a text encoder from Hugging Face333https://huggingface.co/google-bert/bert-base-uncased with a 768-d feature dimension and used no linear layer.
Additional stages
In the additional stage 1.1, we fine-tuned the first-stage audio encoder in the same way as in the fine-tuning experiments in Section IV-B3. In the additional stage 2.1, we pre-trained the second-stage M2D2 with the same setting as in the second stage once again, except for a masking ratio of 0.0 (i.e., no masking) and a limited dataset.
IV-A3 Pre-training datasets
The first stage
In the first stage, the pre-training data consists of AudioSet[17], VGGSound[45], and WavCaps[7], 2,532,215 samples in total. We made it closer to CLAP[6]; they used 4.6 M samples from AudioSet, WavCaps, and many other datasets. We used Auto-ACD[46] for the caption data of AudioSet and VGGSound and supplemented the missing captions with ”The sound of labels” using their labels. Since the curated dataset included data samples for the downstream tasks (AudioCaps[41], Clotho[47], ESC-50[16], US8K[48]) used in the evaluations, we removed all of these data samples. For the audio data, we randomly cropped 6-s segments during the training.
The second stage
For the second stage, based on the first stage dataset, we added the AudioCaps and Clotho training samples for a fair comparison with previous studies, especially WavCaps[7]. As with the first stage, we removed all data samples of ESC-50/US8K from the added AudioCaps/Clotho training samples. The total number of samples was 2,585,227. In this stage, we used 10-s audio data as they were during the training, or we randomly cropped for 10 s if a sample was longer.
Additional stages
For the additional training stages, we used AudioSet containing 2,005,132 training and 20,178 test samples with its labels for fine-tuning in stage 1.1. In stage 2.1, we used WavCaps, AudioCaps, and Clotho, 439,743 samples in total, as in the previous CLAP studies[7][49].
| Environmental sound tasks | Speech tasks | Music tasks | ||||||||||
| AS2M[17] | AS20K[17] | ESC-50[16] | US8K[48] | FSD50K[50] | SPCV2[51] | VC1[18] | VF[18] | CRM-D[52] | GTZAN[53] | NSynth[54] | Surge[55] | |
| Evaluation protocol† | FT | FT / ZS | FT / LE / ZS | LE / ZS | ZS | FT / LE | FT / LE | LE | LE / ZS | LE / ZS | LE / ZS | LE |
| Training samples | 2,005,132 | 21,940 | 5 folds 2000 | 10 folds 8732 | 40,966 | 84,843 | 138,361 | 121,281 | 5155 | 443 | 289,205 | 148,896 |
| Validation samples | - | - | - | 9981 | 6904 | 26,684 | 732 | 197 | 12,678 | 17,160 | ||
| Test samples | 20,178 | 20,178 | 10,231 | 11,005 | 8251 | 28,463 | 1551 | 290 | 4096 | 17,336 | ||
| Classes | 527 | 527 | 50 | 10 | 200 | 35 | 1251 | 6 | 6 | 10 | 11 | 88 |
| Average duration | 10.0 s | 10.0 s | 5.0 s | 4.0 s | 7.6 s | 1.0 s | 8.2 s | 5.8 s | 2.5 s | 30.0 s | 4.0 s | 4.0 s |
| †FT, LE, and ZS stand for fine-tuning, linear evaluation, and zero-shot classification, respectively. | ||||||||||||
under a unified linear evaluation except Pengi, Cacophony, and CoLLAT.
| Audio enc. fine-tuning on AudioSet | Env. sound tasks | Speech tasks | Music tasks | ||||||||
| Model | ESC-50 | US8K | SPCV2 | VC1 | VF | CRM-D | GTZAN | NSynth | Surge | Avg. | |
| (Previous studies: Audio models) | |||||||||||
| ATST-Clip [32] | 94.1 0.6 | 85.8 | 95.1 | 72.0 | 97.6 0.0 | 68.8 1.3 | 78.9 3.5 | 76.2 | 32.8 0.0 | 77.9 | |
| ATST-Frame [32] | 90.9 0.6 | 85.8 | 94.9 | 77.4 | 98.8 0.3 | 72.3 0.7 | 82.9 6.0 | 75.9 | 40.6 0.2 | 79.9 | |
| Cacophony (stage 1)[10] | 87.0 | - | 92.2 | - | - | 69.7 | 83.8 | - | - | - | |
| BEATs [33] | 86.9 1.4 | 84.8 0.1 | 89.4 0.1 | 41.4 0.7 | 94.1 0.3 | 64.7 0.8 | 72.6 4.3 | 75.9 0.2 | 39.3 0.4 | 72.1 | |
| BEATs [33] | (✓) | 95.5 0.3 | 87.6 0.3 | 86.7 0.1 | 37.0 0.2 | 92.5 0.1 | 67.6 1.5 | 84.6 0.5 | 73.1 0.4 | 35.7 0.3 | 73.4 |
| CED [34] | 97.3 0.5 | 87.8 0.2 | 89.0 0.3 | 35.2 0.2 | 94.8 0.1 | 66.1 1.3 | 42.3 15.4 | 75.6 0.5 | 38.9 0.6 | 69.7 | |
| AST[25] | ✓ | 94.60.7 | 85.40.6 | 72.60.0 | 16.50.3 | 81.20.3 | 58.43.1 | 85.10.5 | 72.90.9 | 25.70.1 | 65.8 |
| HTS-AT [26] | ✓ | 95.7 0.7 | 83.8 0.1 | 82.1 0.3 | 18.1 0.4 | 82.3 0.3 | 56.2 0.6 | 85.1 0.5 | 73.3 0.8 | 26.3 0.5 | 67.0 |
| PANNs CNN14[24] | ✓ | 90.70.8 | 81.90.1 | 51.80.3 | 7.80.1 | 75.10.2 | 51.81.3 | 78.73.0 | 65.90.2 | 10.90.3 | 57.2 |
| (Previous studies: CLAP and audio-language models) | |||||||||||
| LAION-CLAP [4] | ✓ | 97.3 0.5 | 86.9 0.5 | 75.9 0.5 | 13.4 0.4 | 80.3 0.2 | 54.6 1.0 | 84.3 2.6 | 72.2 1.1 | 14.8 0.5 | 64.4 |
| CLAP2022 [5] | ✓ | 93.8 0.1 | 84.2 0.7 | 59.0 1.1 | 8.9 0.6 | 75.8 1.3 | 54.4 0.8 | 79.3 | 68.2 0.6 | 8.4 0.7 | 59.1 |
| CLAP2023 [6] | ✓ | 97.7 0.5 | 88.4 0.1 | 86.2 0.8 | 21.1 0.3 | 89.6 0.8 | 62.5 1.8 | 82.3 0.5 | 80.5 0.1 | 27.2 0.5 | 70.6 |
| Pengi [56] | ✓ | 89.15 | - | - | - | - | 50.57 | 80.0 | - | - | - |
| WavCaps [7] | ✓ | 97.2 0.3 | 63.6 0.6 | 73.3 1.7 | 16.9 0.2 | 80.0 1.0 | 58.6 0.7 | 80.2 1.3 | 74.4 0.9 | 21.1 0.2 | 62.8 |
| Cacophony (stage 2)[10] | 97.0 | - | 76.2 | - | - | 59.3 | 85.0 | - | - | - | |
| CoLLAT[9] | ✓ | 97 | 89 | - | - | - | - | - | - | - | - |
| (Baseline: M2D) | |||||||||||
| M2D [23] | 91.3 0.6 | 87.6 0.2 | 96.0 0.1 | 73.4 0.2 | 98.3 0.0 | 73.0 0.7 | 84.1 2.7 | 75.7 0.1 | 42.1 0.2 | 80.2 | |
| M2D[23] | ✓ | 96.7 0.5 | 89.3 0.1 | 93.2 0.2 | 52.7 0.6 | 96.9 0.0 | 69.5 1.8 | 87.6 0.9 | 75.6 0.6 | 40.5 0.4 | 78.0 |
| (Proposed: Audio-language models) | |||||||||||
| M2D2, M2D2-, M2D2+ | 95.2 0.3 | 89.5 0.2 | 95.6 0.1 | 68.6 0.5 | 98.1 0.1 | 73.4 0.7 | 86.6 1.6 | 76.5 0.4 | 42.2 0.3 | 80.6 | |
| M2D2 | ✓ | 97.9 0.5 | 89.7 0.3 | 94.8 0.2 | 56.3 0.5 | 97.1 0.0 | 69.7 0.3 | 86.3 1.8 | 76.7 0.4 | 41.1 0.3 | 78.8 |
| (Ablation using a non-LLM text encoder and AudioSet only as audio data. The underlined results outperform M2D2-) | |||||||||||
| M2D20-(M2D-CLAP[11]) | 96.3 0.3 | 88.8 0.6 | 95.8 0.3 | 70.3 0.4 | 98.3 0.1 | 73.4 0.2 | 84.1 1.5 | 78.0 0.5 | 42.4 0.6 | 80.8 | |
| Results quoted from corresponding papers when they are better than ours or unavailable in our test. | |||||||||||
| † Results using the HEAR benchmark[57] in the settings similar to those for the linear evaluation. | |||||||||||
IV-A4 Evaluation tasks
Table I lists the downstream tasks, which include environmental sound tasks (AudioSet[17], ESC-50[16], UrbanSound8K[48] (US8K), and FSD50K), speech tasks (Speech Commands V2[51] (SPCV2), VoxCeleb1[18] (VC1), VoxForge[58] (VF), CREMA-D[52] (CRM-D)), and music tasks (GTZAN[53], NSynth[54], and Pitch Audio Dataset (Surge synthesizer)[55] (Surge)). AudioSet has two settings: AS2M with all 2M training samples and AS20K with 21K samples from the balanced train segments, and Surge is a non-semantic pitch audio classification of 88 MIDI notes.
IV-B Evaluation: Audio Features
We evaluated the effectiveness of the pre-trained audio features under the frozen feature and fine-tuning settings on general tasks and the frozen feature setting on music tasks. We followed [15][23] for the frozen feature (linear evaluation) and fine-tuning experimental setup using a unified evaluation platform, EVAR444https://github.com/nttcslab/eval-audio-repr, on general tasks, and used MARBLE benchmark[59] for the music task evaluation.
IV-B1 Frozen audio features on general audio tasks
Experimental setup
We followed the standard linear evaluation procedure[15][23]. Specifically, the evaluation pipeline first converts downstream task samples into features using a frozen pre-trained model, trains a linear layer with task labels, and then conducts tests for the trained linear layer. We used the validation set for early stopping with patience of 20 epochs and trained the linear layer for up to 200 epochs using the Adam optimizer and a learning rate of 0.00003.
Results: Frozen audio features on general audio tasks
The results of the linear evaluation shown in Table II indicate that supervised learning and CLAP models perform poorly on some tasks, whereas the M2D2/M2D variants generally perform well. For example, supervised learning models AST, HTS-AT, and PANNs show 7.8 to 18.1 on VC1, and CLAP models such as LAION-CLAP show 8.4 to 27.2 on Surge, whereas the M2D variants show 52.7 to 73.4 on VC1 and 40.5 to 42.4 on Surge; we also observe these differences in other speech tasks. Note that the CLAP models use supervised learning models as their audio encoders; hence, the results indicate that applying audio features learned using labels/captions to tasks outside the coverage of these signals is challenging. For example, a training signal of ”Male speech, man speaking” does not cover who is speaking on VC1 (speaker identification).
BEATs and CED, pre-trained with distillation of the previous training cycle models, show a similar trend to supervised learning models. CED shows 35.2 on VC1, and BEATs shows 35.7 on Surge, for example. These models have distilled the pre-trained models on AudioSet, making them focused more on AudioSet data distribution. The results also indicate that their features become difficult to apply for tasks with different data distributions.
In contrast, M2D2/M2D variants and an SSL model ATST perform well in all tasks. M2D2 improves performance on ESC-50 at 94.9 and US8K at 89.1 compared to M2D (note that M2D2-/M2D2+ share the same audio encoder with M2D2). However, it exhibits slight performance degradation in the speech task, especially VC1, indicating a trade-off between M2D and CLAP’s training signal.
Notably, M2D2 (M2D2 encoder fine-tuned to AudioSet) does not degrade performance severely in speech tasks while improving performance on ESC-50 and US8K. It shows SOTA performance of 97.9 on ESC-50 and 89.7 on US8K. However, it degrades performance to 56.3 on VC1 from M2D2’s 68.6, while it is still in the good performance range compared to previous models. Similarly, M2D, a fine-tuned version of the original M2D, improves performance on ESC-50/US8K and degrades on others slightly more than M2D2 but not as severely as the supervised learning models. These results show that the M2D pre-training helps learn general-purpose representations that do not severely compromise performance against subsequent fine-tuning.
In addition, M2D2, which is pre-trained only on AudioSet using GTE-Base[60] (non-LLM-based sentence embedding model) as a text encoder, performs better than M2D2 in most tasks, indicating that the pre-training setting, especially relying solely on AudioSet is beneficial for performing well with frozen features. However, as shown in Section IV-C, its CLAP features severely underperform M2D2.
In summary, the experiments validated the effectiveness of M2D2 and the variants as a frozen audio extractor on general audio tasks.
| MTT | GTZAN | EMO | NSynth | VocalSet | MTG | ||||||||||||
| Tagging | Genre | Emotion | Inst. | Pitch | Tech | Singer | Instrument | MoodTheme | Genre | Top50 | |||||||
| Model | ROC | AP | Acc | R2V | R2A | Acc | Acc | Acc | Acc | ROC | AP | ROC | AP | ROC | AP | ROC | AP |
| (Previous studies: Audio and audio-language models) | |||||||||||||||||
| ATST-Frame[32] | 91.2 | 39.7 | 79.0 | 60.3 | 76.7 | 77.0 | 84.2 | 76.8 | 88.7 | 78.5 | 21.0 | 77.0 | 15.1 | 87.0 | 19.5 | 83.0 | 29.7 |
| BEATs[33] | 91.6 | 40.7 | 84.6 | 55.1 | 70.9 | 78.1 | 84.2 | 74.8 | 84.8 | N/A (unable to handle long audio) | |||||||
| CLAP2023[6] | 90.0 | 38.6 | 79.0 | 47.6 | 68.1 | 81.0 | 68.1 | 70.7 | 73.7 | 76.2 | 18.5 | 75.4 | 12.9 | 83.9 | 15.6 | 80.8 | 26.2 |
| (Previous studies: SOTA music audio models) | |||||||||||||||||
| Jukebox-5B[61] | 91.4 | 40.6 | 77.9 | 57.0 | 73.0 | 70.4 | 91.6 | 76.7 | 82.6 | 78.5 | 22.0 | 77.6 | 15.3 | 88.0 | 20.5 | 83.7 | 30.6 |
| MULE[62] | 91.2 | 40.1 | 75.5 | 60.7 | 73.1 | 74.6 | 88.5 | 75.5 | 87.5 | 76.6 | 19.2 | 78.0 | 15.4 | 88.0 | 20.4 | 83.7 | 30.6 |
| MAP-MERT-v1-95M[63] | 91.0 | 39.3 | 74.8 | 55.5 | 76.3 | 70.7 | 92.6 | 74.2 | 83.7 | 77.5 | 19.4 | 76.4 | 13.4 | 87.1 | 18.8 | 83.0 | 29.0 |
| MAP-MERT-v1-330M[63] | 91.1 | 39.5 | 77.6 | 59.0 | 75.8 | 72.6 | 94.4 | 76.9 | 87.1 | 78.1 | 19.8 | 76.5 | 14.0 | 86.7 | 18.6 | 83.4 | 29.9 |
| (Baseline: M2D) | |||||||||||||||||
| M2D | 91.4 | 40.5 | 81.5 | 59.4 | 76.1 | 78.2 | 89.3 | 76.1 | 92.0 | 79.2 | 21.1 | 77.6 | 15.7 | 87.6 | 20.1 | 83.4 | 29.8 |
| M2D | 91.5 | 40.9 | 86.1 | 59.0 | 73.0 | 78.7 | 88.2 | 76.8 | 91.3 | 79.7 | 23.3 | 77.8 | 15.6 | 87.5 | 20.3 | 83.4 | 30.6 |
| (Proposed: Audio-language models) | |||||||||||||||||
| M2D2, M2D2-, M2D2+ | 91.5 | 40.8 | 82.6 | 59.3 | 76.7 | 77.4 | 89.1 | 77.4 | 91.8 | 79.3 | 21.7 | 78.0 | 15.9 | 87.9 | 20.8 | 83.7 | 30.3 |
| M2D2 | 91.8 | 41.6 | 85.9 | 58.2 | 76.2 | 79.7 | 89.5 | 78.9 | 92.7 | 79.7 | 22.2 | 78.8 | 15.8 | 88.3 | 20.9 | 84.0 | 30.9 |
| (Ablation using a non-LLM text encoder and AudioSet only as audio data. The underlined results outperform M2D2-) | |||||||||||||||||
| M2D20-(M2D-CLAP[11]) | 91.6 | 41.0 | 84.3 | 61.9 | 77.4 | 80.6 | 88.8 | 78.1 | 92.3 | 79.0 | 21.6 | 78.2 | 16.0 | 88.0 | 20.8 | 83.8 | 30.4 |
IV-B2 Frozen audio features on music tasks
To further assess the frozen feature, we evaluated models on the music audio tasks with various music audio signals.
Experimental setup
We used the MARBLE benchmark[59], which features diverse music audio tasks for music tracks, instrument sounds, and vocals. We used the tasks for which we could reproduce the results in the paper with the released code555https://github.com/a43992899/MARBLE-Benchmark, including music tagging and genre classification on MagnaTagATune (MTT)[64], GTZAN[53], and MTG-Jamendo (MTG)[65], emotion detection on Emomusic[66] and MTG, pitch and instrument (inst) classification on NSynth[54] and MTG, and vocal technique detection (tech) and singer identification on VocalSet[67].
The evaluation metrics are the area under the ROC curve (ROC) and the average precision (AP), except for Emomusic, which uses the determination coefficient between the model’s regression results and human annotations of valence (R2V) and arousal (R2A)[66]. The results are the averages of five tests with different random seeds.
Results: Frozen audio features on music tasks
The results of the music task benchmark in Table III show that general audio models perform comparably to the SOTA music audio models. They even outperform music models on GTZAN and NSynth (Inst) while underperforming them in pitch classification NSynth (Pitch). While music models are pre-trained on large-scale music databases, general models are also pre-trained using large-scale music data because more than half (1M) of the AudioSet samples are tagged as ”music”[17]. In addition, most of the music and general models share a similar SSL-based training scheme. These facts would explain the comparable performance of general models on the music-specific benchmark. However, the CLAP model, CLAP2023, using a supervised learning audio backbone, underperforms other general and music models except for NSynth Inst666NSynth is contained in the CLAP2023 training data., suggesting the need for SSL to generalize in music tasks. In contrast, the SSL general audio models perform well in music tasks, indicating that pre-training on AudioSet with SOTA SSL methods can generalize well.
The models pre-trained using both SSL and supervised learning of AudioSet (M2D, M2D2, and BEATs) show stronger performance, and M2D2 demonstrates new SOTA results over the SOTA music audio models. In addition, as in the frozen audio feature evaluation on general tasks, M2D2 performs better than M2D2 in most music tasks, also indicating that pre-training solely on AudioSet is beneficial for performing well with frozen features. To summarize, the experiment confirmed that M2D2 models perform well and show SOTA performance in diverse music tasks.
IV-B3 Fine-tuning
Experimental setup
We followed M2D[23], ATST[32], and Audio-MAE[31] for the fine-tuning procedure. Specifically, a linear classifier was added on top of the pre-trained model to train the entire network. We used SpecAugment[68], Mixup[69, 15], and Random Resize Crop (RRC)[15] for data augmentation, along with Structured Patchout[70] that randomly masks patches during training. We used SGD, LARS, AdamW optimizers, and cosine annealing[71] learning rate scheduling. We used the hyperparameters for each task as listed in Table IV. For Structured Patchout, we used 768-d features calculated by averaging all ViT output features because Eq. (2) is not applicable to the masked patches. We used the tasks commonly used in previous studies, namely ESC-50, SPCV2, and VC1, the same as in the linear evaluation, plus AS2M and AS20K.
| Parameter | AS2M | AS20K | ESC-50 | SPCV2 | VC1 |
| Learning rate | 2.0 | 0.5 | 0.5 | 0.5 | 0.0005 |
| Batch size | 64 | 64 | 128 | 128 | 64 |
| Optimizer | LARS | SGD | SGD | SGD | AdamW |
| Mixup ratio | 0.5 | 0.3 | 0.0 | 0.3 | 0.0 |
| Random resize crop | - | ✓ | ✓ | ✓ | - |
| SpecAugment†[68] | 30/192 | 30/192 | 15/48 | 30/48 | 30/48 |
| Training epochs (total) | 70 | 200 | 200 | 200 | 50 |
| Training epochs (warm-up) | 15 | 5 | 5 | 5 | 5 |
| Structured Patchout[70] ratio | 0.5 | 0.5 | 0.5 | 0.5 | 0.0 |
| Adjust positional encoding | ✓ | ✓ | - | - | ✓ |
| Freeze embedding layer[72] | - | - | ✓ | - | - |
| † Frequency/time masking parameters. | |||||
| AS2M | AS20K | ESC-50 | SPCV2 | VC1 | |
| Model | mAP | mAP | acc(%) | acc(%) | acc(%) |
| (Previous studies: Audio models) | |||||
| BEATs [33] | 48.0 | 38.3 | 95.6 | 98.3 | - |
| BEATs [33] | 48.6 | 41.8 | 98.1 | 98.1 | - |
| ATST-Clip [32] | 45.2 | 37.9 | - | 98.0 | 95.5 |
| ATST-Frame [32] | 48.0 | 39.0 | - | 98.1 | 97.3 |
| HTS-AT [26] | 47.1 | - | 97.0 | 98.0 | - |
| DTF-AT Full TF[73] | 48.6 | - | 97.5 | 97.68 | - |
| (Previous studies: Audio-language models) | |||||
| AudioCLIP [2] | - | - | 97.15 | - | - |
| Wav2CLIP [3] | - | - | 85.95 | - | - |
| CLAP2022 [5] | - | - | 96.7 | 96.8 | - |
| (Baseline: M2D) | |||||
| M2D | 47.9 0.0 | 38.6 0.1 | 96.0 0.2 | 98.4 0.1 | 96.3 0.2 |
| (Proposed: Audio-Language models) | |||||
| M2D2, M2D2-, M2D2+ | 49.0 0.1 | 42.1 0.0 | 97.8 0.1 | 98.4 0.0 | 95.5 0.3 |
| M2D2 | - | 46.7 0.2 | 98.5 0.0 | 98.4 0.1 | 94.0 0.2 |
| (Ablation using a non-LLM text encoder and AudioSet only as audio data) | |||||
| M2D20-(M2D-CLAP[11]) | 48.5 0.1 | 41.8 0.2 | 97.4 0.2 | 98.3 0.1 | 95.5 0.2 |
| (Reference: Models using ensemble of audio models or distillation from large models) | |||||
| CED [34] | 50.0 | 44.0 | 96.65 | - | - |
| ATST-C2F [32] | 49.7 | 40.5 | - | 98.4 | 97.5 |
| SupMAM-CLAP [38] | 48.5 | 38.6 | 97.6 | 98.7 | - |
| DyMN-L[74] | 49.0 | - | 97.4 | - | - |
Results: Fine-tuning
The experimental results in Table V show that M2D2 variants further improve the performance of the baseline M2D. Unlike the linear evaluation that evaluated all models except a few, we evaluated only our models in fine-tuning because reproducing the best results of the previous methods is quite demanding.
M2D2 improves the performance from M2D, which shows SOTA results of 49.0 for AS2M and top-level results of 42.1 for AS20K and 97.8 for ESC-50. On the other hand, VC1 performance is 95.5, which is lower than the original M2D’s performance of 96.3, suggesting a trade-off for learning CLAP. M2D2 with an audio encoder fine-tuned to AudioSet exhibits an even stronger trend, showing SOTA results of 46.7 for AS20K and 98.5 for ESC-50 while 94.0 for VC1, a -1.5 pp degradation compared to M2D2.
The experiment confirmed that M2D2 achieves audio features that exhibit SOTA performance. Notably, it shows top performance with a single pre-training, unlike other SOTA models that require multi-iteration/model pre-training. In addition, the performance boost of M2D2 from M2D20- (M2D-CLAP) suggests the benefit of using a powerful LLM-based sentence encoder.
IV-C Evaluation: CLAP Features
We evaluated the effectiveness of the CLAP features in zero-shot classification, audio-text retrieval, and audio captioning experiments.
IV-C1 Zero-shot Classification
We conducted experiments to evaluate the zero-shot classification performance of CLAP features. Under the zero-shot classification setting, a model infers the class label of audio by picking a class with the highest similarity between a CLAP audio feature and CLAP text feature of each class.
Experimental setup
We chose the tasks found in previous papers, which include ESC-50 (ESC), US8K, CREMA-D (CRD), GTZAN (GTZ), NSynth (NS), AudioSet (AS), and FSD50K (FSD). We used AudioSet and FSD50K as references because M2D2 was pre-trained on AudioSet and part of FSD50K in WavCaps. We conducted experiments using a standard procedure to infer a class label of audio with the closest cosine distance between the features of each test sample and the label’s caption, and we obtained accuracies from the inference results. Table VI summarizes the rules for converting task labels to the captions. We used the evaluation platform EVAR, as in Section IV-B, to evaluate M2D2 and pervious models.
Results: Zerio-shot-classification
Table VII shows the experimental results. We compare results based on a strict zero-shot classification setting as in [7]; we only include experimental results that do not use data from downstream tasks during training in our comparisons and gray out all other results as references.
Results show that M2D2 variants achieve top-level performance on all tasks. Among the variants, M2D2- outperforms previous models with a SOTA result of 79.31 for GTZAN, while the second-stage pre-training degrades performance, such as 72.41 with M2D2. M2D2+ brings the ESC-50 and US8K results to the top level, indicating that CLAP training with the limited combination of WavCaps, AudioCaps, and Clotho is advantageous for these tasks compared to M2D2 and M2D2- that use large-scale datasets. Furthermore, M2D2- significantly improves performance on these tasks from M2D20-, indicating the benefit of using an LLM-based text encoder. On the other hand, M2D2 improves CREMA-D closer to the top, CLAP2023.
In summary, we validated that M2D2 achieves top-level performance in the zero-shot classification setting.
| Task | Rule |
| AS & FSD | ”labels can be heard” |
| ESC50 & US8K | ”label can be heard” |
| CREMA-D | ”(angry person talking someone talking in disgust |
| someone talking with a sense of fear | |
| someone talking happily and joyfully | |
| someone talking calmly someone talking sadly) can be heard” | |
| GTZAN | ”label music can be heard” |
| NSynth | ”the musical instrument sound of label can be heard” |
| AS | FSD | ESC | US8K | CRD | GTZ | NS | |
| Model | mAP | mAP | acc(%) | acc(%) | acc(%) | acc(%) | acc(%) |
| (Previous studies) | |||||||
| AudioCLIP [2] | - | - | 69.40 | 68.78 | - | - | - |
| Wav2CLIP [3] | - | 3.02 | 41.4 | 40.44 | - | - | - |
| WavCaps [7] | 19.60 | 52.96 | 94.8 | 81.42 | 19.86 | 45.52 | 27.66 |
| LAION-CLAP [4] | - | 45.85 | 91.0 | 77.0 | 23.08 | 47.24 | 35.28 |
| CLAP2022 [5] | 5.8 | 30.24 | 82.6 | 75.29 | 22.76 | 28.97 | 21.44 |
| CLAP2023 [6] | 10.2 | 48.5 | 93.90 | 82.3 | 30.0 | 58.4 | 58.08 |
| MGA-CLAP+[35] | 23.0 | 54.5 | 94.9 | 83.7 | - | - | - |
| Cacophony[10] | - | - | 93.4 | 77.1 | - | - | - |
| Pengi [56] | 16.35 | 46.76 | 91.95 | 71.85 | 18.46 | 35.25 | 50.07 |
| LTU[75]/-AS[76] | 18.7 | 46.3 | 83.1 | - | - | 50.3 | - |
| CoLLAT[9] | 9 | 19 | 84 | 77 | - | - | - |
| JMLA [77] | - | - | - | - | - | 64.82 | - |
| (Proposed models) | |||||||
| M2D2- | 23.15 | 43.43 | 90.07 | 76.52 | 17.41 | 79.31 | 31.76 |
| M2D2 | 28.27 | 46.76 | 93.67 | 81.92 | 23.50 | 72.41 | 24.57 |
| M2D2+ | 27.13 | 42.03 | 94.55 | 82.90 | 21.73 | 70.86 | 29.52 |
| M2D2 | 30.05 | 50.04 | 94.30 | 82.30 | 28.56 | 70.69 | 30.57 |
| (Ablation using a non-LLM text encoder and AudioSet only as audio data) | |||||||
| M2D20-(M2D-CLAP) | 27.24 | 40.82 | 75.45 | 72.40 | 17.73 | 75.17 | 23.39 |
| (Reference studies that employ additional techniques on top of existing CLAP models.) | |||||||
| Proto-LC [78] | - | 52 | 96 | 73 | - | - | - |
| CD [79] | 19.98 | 50.74 | 96.35 | 90.17 | - | - | - |
| Results quoted from papers when they are better than our reproduction or test unavailable. | |||||||
| Final CLAP training dataset | Audio encoder AudioSet fine-tuning | AudioCaps | Clotho | |||||||||||||||
| Text-to-Audio | Audio-to-Text | Text-to-Audio | Audio-to-Text | |||||||||||||||
| CLAP model | R@1 | R@5 | R@10 | mAP@10 | R@1 | R@5 | R@10 | mAP@10 | R@1 | R@5 | R@10 | mAP@10 | R@1 | R@5 | R@10 | mAP@10 | ||
| (Previous studies) | ||||||||||||||||||
| HTSAT-BERT-ZS[7] | WC | ✓ | 28.6 | 61.1 | 75.8 | - | 40.2 | 69.4 | 80.3 | - | 16.5 | 38.8 | 50.9 | - | 20.0 | 43.3 | 56.6 | - |
| HTSAT-BERT-PT[7] | WC+A+C | ✓ | 39.7 | 74.5 | 86.1 | - | 51.7 | 82.3 | 90.6 | - | 19.5 | 45.2 | 58.2 | - | 23.4 | 50.9 | 63.4 | - |
| HTSAT-BERT-FT[7] | A or C | ✓ | 42.2 | 76.5 | 87.1 | - | 54.6 | 85.2 | 92.4 | - | 19.7 | 45.7 | 59.4 | - | 26.9 | 52.6 | 64.90 | - |
| CNN14-BERT-FT[7] | A or C | ✓ | 35.1 | 70.0 | 82.1 | - | 45.7 | 76.1 | 87.7 | - | 21.5 | 47.9 | 61.9 | - | 27.1 | 52.7 | 66.3 | - |
| LAION-CLAP[4] | LS+LA+A+C | ✓ | 35.1 | 71.9 | 83.7 | - | 44.2 | 80.8 | 90.3 | - | 16.9 | 41.6 | 54.4 | - | 24.4 | 49.3 | 65.7 | - |
| CLAP2023[6] | LS+WC+A+C | ✓ | 35.6 | - | - | 51.0 | 42.5 | - | - | 31.9 | 15.7 | - | - | 25.7 | 22.9 | - | - | 15.5 |
| MGA-CLAP∗[35] | WC+A+C | ✓ | 42.2 | 74.9 | - | - | 53.7 | 84.3 | - | - | 20.8 | 45.0 | - | - | 26.5 | 54.1 | - | - |
| Cacophony[10] | LS+A+C | 41.0 | 75.3 | 86.4 | - | 55.3 | 83.6 | 92.4 | - | 20.2 | 45.9 | 58.8 | - | 26.5 | 54.1 | 67.3 | - | |
| FLAP[8] | LA+A+C | 41.5 | 75.5 | 86.0 | - | 53.0 | 84.1 | 92.6 | - | 20.3 | 46.5 | 58.8 | - | 25.5 | 53.4 | 67.9 | - | |
| (Proposed models) | ||||||||||||||||||
| M2D2 | LS+WC | 27.34 | 58.14 | 71.45 | 40.14 | 36.05 | 65.31 | 77.43 | 23.95 | 15.92 | 37.88 | 49.26 | 25.37 | 18.09 | 39.90 | 52.54 | 12.39 | |
| M2D2 | LS+WC+A+C | 37.24 | 72.52 | 84.54 | 52.00 | 52.56 | 82.34 | 90.60 | 38.74 | 18.09 | 42.53 | 55.52 | 28.51 | 24.88 | 48.13 | 60.48 | 15.53 | |
| M2D2+ | WC+A+C | 40.54 | 74.52 | 86.92 | 55.09 | 57.99 | 83.91 | 91.43 | 41.72 | 19.43 | 45.84 | 58.64 | 30.55 | 26.32 | 51.20 | 65.07 | 17.65 | |
| M2D2 | WC+A+C | ✓ | 41.96 | 77.07 | 88.57 | 56.74 | 59.25 | 84.74 | 92.79 | 43.24 | 20.08 | 45.99 | 59.48 | 31.21 | 24.98 | 51.67 | 64.59 | 17.35 |
| (Ablations using a non-LLM text encoder. For the audio data, M2D20-(M2D-CLAP) used AudioSet only while +WC used AudioSet and WavCaps) | ||||||||||||||||||
| M2D20-(M2D-CLAP) | LS | 24.97 | 58.58 | 72.50 | 39.00 | 25.50 | 58.93 | 76.07 | 18.59 | 10.24 | 26.49 | 37.07 | 17.33 | 11.00 | 25.93 | 35.31 | 7.20 | |
| +WC | LS+WC | 26.06 | 57.89 | 72.64 | 39.79 | 28.63 | 61.65 | 75.97 | 19.69 | 14.81 | 35.44 | 47.00 | 23.78 | 15.41 | 33.78 | 47.27 | 10.42 | |
IV-C2 Audio-Text Retrieval
We evaluated the effectiveness of the CLAP features in the audio-text retrieval task setting. This task involves searching for text from audio or audio from text, so unlike zero-shot classification, it requires fine-grained comparisons to find pairs at the data instance level[80]. Therefore, the more accurately the features represent the audio and text, the better the task performance.
Experimental setup
As in previous studies[7][80], we evaluated the CLAP features using the test split samples of AudioCaps[41] and Clotho v2[47]. In addition to the standard metrics recall at rank k (R@k), R@1, R@5, and R@10, we also report mAP@10, the primary metric for the DCASE 2024 Challenge777https://dcase.community/challenge2024/task-language-based-audio-retrieval Task 8.
Results: Audio-text retrieval
Table VIII shows that M2D2- improves overall performance compared to M2D20- and M2D20-+WC, which is pre-trained on the same dataset as M2D2- to make the difference only in the text encoder and audio projector used. M2D2- improves performance on most metrics, e.g., AudioCaps Text-to-Audio R@1 performance of 27.34 compared to M2D20-+WC’s 26.06. The results confirm the effectiveness of the M2D2- updates from M2D20- (M2D-CLAP) using an LLM-based text (sentence) encoder and transformer-based audio projector.
M2D2 further improves overall performance compared to M2D2-, by pre-training the text encoder in the second stage. For example, the AudioCaps R@1 performance improved from M2D2-’s 27.34 to M2D2’s 37.24. The results demonstrate the importance of training the text encoder to achieve fine-grained alignment of the paired features.
The results in Table VIII indicate that the combination of datasets significantly impacts performance. The “Final CLAP training dataset” column in the table lists the datasets used in the model’s last training just before the evaluation. LAION-CLAP, CLAP2023, and M2D2, which used a large-scale dataset (denoted as ”LS” in the table) just before the evaluation, perform poorly compared to those without one. Their AudioCaps Text-to-Audio R@1 performance is about 35 to 37, which is inferior to the 39.7 of HTSAT-BERT-PT without using large-scale datasets, for example. These results indicate that a limited combination of WavCaps, AudioCaps, and Clotho is advantageous for evaluation in the benchmark. As an exception, Cacophony, which developed its data curation strategy, performs well with large-scale datasets.
M2D2+ and M2D2 achieve top performance when additionally trained under conditions advantageous to the benchmark as in the previous studies[4][6][7]. M2D2+ outperforms or becomes comparable to HTSAT-BERT-PT after pre-training on the same data (WavCaps, AudioCaps, and Clotho) as HTSAT-BERT-PT in additional stage 2.1. M2D2 even outperforms top performers HTSAT-BERT-FT, Cacophony, and FLAP in AudioCaps after fine-tuning its audio encoder to AudioSet in the additional stage 1.1.
In summary, the experiments validated the effectiveness of M2D2. M2D2 performs comparably to previous studies, and it further improves performance to the top level through additional stages for a fair comparison to previous studies.
| Audio enc. AudioSet fine-tuning | AudioCaps (subset of AudioSet) | Clotho (out of AudioSet) | |||||||||
| CLAP model | METEOR | CIDEr | SPICE | SPIDEr | FENSE | METEOR | CIDEr | SPICE | SPIDEr | FENSE | |
| (Baselines reproduced on the same codebase) | |||||||||||
| LAION-CLAP[4] | ✓ | 23.8 0.2 | 71.7 0.9 | 17.6 0.2 | 44.6 0.5 | 64.0 0.3 | 17.4 0.1 | 41.4 0.3 | 12.1 0.1 | 26.8 0.2 | 48.6 0.2 |
| CLAP2023[6] | ✓ | 22.4 0.5 | 65.2 2.1 | 16.3 0.3 | 40.7 1.2 | 60.5 0.8 | 17.5 0.1 | 42.5 0.5 | 12.3 0.1 | 27.4 0.3 | 49.2 0.2 |
| (Proposed models) | |||||||||||
| M2D2- | - | 23.3 0.3 | 68.3 1.5 | 16.9 0.2 | 42.6 0.8 | 61.9 0.5 | 17.4 0.1 | 41.2 0.3 | 12.2 0.1 | 26.7 0.2 | 48.9 0.2 |
| M2D2 | - | 23.7 0.3 | 69.6 1.3 | 17.2 0.3 | 43.4 0.7 | 62.5 0.4 | 17.8 0.1 | 43.5 0.4 | 12.4 0.1 | 28.0 0.2 | 49.9 0.2 |
| M2D2+ | - | 23.6 0.3 | 69.4 1.4 | 17.1 0.3 | 43.2 0.8 | 62.3 0.5 | 17.8 0.1 | 43.2 0.4 | 12.5 0.1 | 27.8 0.3 | 49.7 0.2 |
| M2D2 | ✓ | 24.3 0.4 | 72.4 1.8 | 17.6 0.3 | 45.0 1.0 | 63.9 0.4 | 17.7 0.1 | 43.1 0.7 | 12.5 0.2 | 27.8 0.4 | 49.6 0.4 |
| (Reference results from non-EnCLAP models) | |||||||||||
| CNN14-BART[7] | ✓ | 24.7 | 75.6 | 17.9 | 46.8 | - | 18.5 | 48.8 | 13.3 | 31.0 | - |
| HTSAT-BART[7] | ✓ | 25.0 | 78.7 | 18.2 | 48.5 | - | 18.4 | 46.2 | 13.3 | 29.7 | - |
| CLAP2023[6] | ✓ | - | - | - | 45.5 | - | - | - | - | 27.1 | - |
| Cacophony[10] | 23.6 | 72.8 | 16.8 | 44.8 | - | 15.3 | 34.2 | 10.6 | 22.4 | - | |
IV-C3 Audio Captioning
We evaluated the effectiveness of the CLAP audio feature on the audio captioning task using a baseline system that relies on a CLAP model to extract an audio semantic feature to generate audio captions. In the baseline system, the CLAP audio features provide pivotal information to represent the input audio semantics; thus, the better these features represent the audio semantics, the better the output captions.
Experimental setup
We used the official codebase of a SOTA method, EnCLAP888https://github.com/jaeyeonkim99/EnCLAP[12], and replaced the CLAP used in EnCLAP, thereby evaluating only the differences in CLAP under a unified environment. EnCLAP uses CLAP audio features to represent the semantics of the audio clip and EnCodec[81] tokens as supplemental time series information to generate captions using BART[82]. We replaced the CLAP only in the codebase during the evaluation.
We conducted evaluations on AudioCaps and Clotho and used a learning rate of for AudioCaps and for Clotho. We used the same settings for other hyperparameters as in EnCLAP, such as training epochs of 15. We tested the top-five checkpoints of SPIDEr validation performance on the test set and averaged the results to obtain the final result. The evaluation metrics are METEOR[83], CIDEr[84], SPICE[85], SPIDEr[86], and FENSE[87].
Results: Audio captioning
Experimental results in Table IX show that M2D2 is comparable to LAION-CLAP on AudioCaps, while M2D2, M2D2+, and M2D2 outperform LAION-CLAP on Clotho. We also evaluated CLAP2023; however, it underperformed LAION-CLAP. Note that the experiments used the same EnCLAP codebase; the only difference is the CLAP audio feature used by the EnCLAP.
The AudioCaps results demonstrate that fine-tuning the audio encoder to AudioSet is vital for good performance; M2D2 performs comparably to LAION-CLAP while M2D2+ does not. On the other hand, the Clotho results suggest that fine-tuning to AudioSet does not offer general-purpose performance; M2D2 and M2D2+ perform comparably to M2D2, and they all outperform LAION-CLAP. Notably, M2D2 performs slightly better than M2D2+ and M2D2 on Clotho, demonstrating that it learns generalizable CLAP features without the need for benchmark techniques.
In summary, the experimental results confirm that the M2D2 models achieve CLAP features that also perform well in audio captioning.
IV-D M2D2 Analysis
IV-D1 Taming the CLAP objective in the first stage
While training with the CLAP objective provides a powerful training signal for the semantic features, it is challenging to balance it with M2D masked prediction learning, which trains effective general-purpose representations. We investigated the impact of CLAP loss weight adjustment and audio projector design on the performance of the learned representations.
CLAP loss weight ablations
We investigated the effect of CLAP loss weight in Eq. (5), which directly adjusts the effect of the CLAP objective. Fig. 4 shows the average results of linear evaluation on general tasks across training epochs for (a) all tasks, (b) ESC-50/US8K tasks for which the CLAP objective is effective, and (c) all other tasks when the weight is varied.
When the weight is at 0.1, (b) ESC-50/US8K performance matures quickly and degrades as training progresses, while (c) performance for other tasks does not improve. When the weight is at 0.001, (b) ESC-50/US8K performance improves slowly, suggesting the weight is too low. In contrast, with the weight at 0.01, performance for all tasks matures as training progresses.
The ESC-50/US8K performance’s maturing quickly with high weight is similar to the observation that the CLAP training matures early in the previous study[6]. The observations of all cases suggest that combining SSL and the CLAP objectives while suppressing CLAP is essential for successful pre-training.
Audio projector ablations
The audio projector, which aggregates the audio features of each token into a single feature, impacts the effectiveness of training through CLAP loss calculation. Fig. 5 shows the average results of linear evaluation across training epochs for the ablations of audio projectors: M2D2’s one-block transformer encoder, the two-block encoder (one additional block), and one-hidden-layer MLP used in M2D-CLAP (M2D20-). When using MLP as a projector, we averaged all the audio encoder output features and fed the averaged feature to it.
The results in Fig. 5(a) show that MLP exhibits the best results up to 200 epochs but stalls afterward. In the final 300 epochs, one block transformer of M2D2 outperforms the others and obtains the best results. The MLP results in Fig. 5(b) show a trend similar to the ESC-50/US8K results when the CLAP loss weight is high () in Fig. 4(b), suggesting that the use of a simple MLP augments the contribution of the CLAP objective during training and leads to early maturation, as reported in the previous study[6].
The two-block transformer underperformed the one-block transformer throughout the training process, indicating that using a projector with adequate network capacity is also vital for successful pre-training.
IV-D2 Text encoder ablations
We assessed the performance impact on the type of text encoder used in the first stage. We compared the LLM-based model NV-Embed-v2[21][42] (7.85B parameters) used in M2D2 against the GTE-Base[60] (109M parameters) used in M2D-CLAP (M2D20-) and GTE-Qwen2-7B-instruct999https://huggingface.co/Alibaba-NLP/gte-Qwen2-7B-instruct (7.61B parameters)[60], which ranked higher than NV-Embed-v2 on the MTLB leaderboard. The experiments replaced the text encoder in M2D2. We also assessed the performance of the second-stage text encoder when using GTE-Base instead of BERT.
The linear evaluation results on the general audio tasks in Table X and the audio-text retrieval of the CLAP features in Table XI show that using NV-Embed-v2’s LLM-based sentence features achieves the best performance with both features. With a non-LLM-based GTE-Base, M2D2 learns competitive audio features while underperforming second-stage CLAP features compared to LLM-based encoders. The GTE-Qwen2-7B-instruct’s results are comparable to (while slightly worse than) NV-Embed-v2’s, confirming the effectiveness of employing LLM-based text encoders.
| Text encoder | ESC | US8K | SPC | VC1 | VF | CRD | GTZ | NS | SUR | Avg. |
| NV-Embed-v2 | 95.2 | 89.5 | 95.6 | 68.6 | 98.1 | 73.4 | 86.6 | 76.5 | 42.2 | 80.6 |
| GTE-Qwen2-7B | 95.4 | 89.1 | 95.9 | 67.7 | 98.1 | 73.6 | 85.2 | 75.0 | 42.4 | 80.3 |
| GTE-Base | 95.3 | 89.2 | 95.6 | 66.2 | 97.7 | 74.7 | 84.3 | 75.2 | 42.2 | 80.0 |
| Text encoder used in the first stage | AudioCaps | Clotho | ||||||
| T-to-A | A-to-T | T-to-A | A-to-T | |||||
| R@1 | mAP @10 | R@1 | mAP @10 | R@1 | mAP @10 | R@1 | mAP @10 | |
| (The first-stage pre-training) | ||||||||
| NV-Embed-v2 | 27.34 | 40.14 | 36.05 | 23.95 | 15.92 | 25.37 | 18.09 | 12.39 |
| GTE-Qwen2-7B | 21.88 | 34.35 | 31.45 | 20.05 | 14.37 | 22.88 | 16.08 | 10.31 |
| GTE-Base | 26.96 | 40.45 | 32.18 | 21.95 | 14.93 | 23.85 | 17.03 | 11.34 |
| (The second-stage pre-training using BERT as a text encoder) | ||||||||
| NV-Embed-v2 | 37.24 | 52.00 | 52.56 | 38.74 | 18.09 | 28.51 | 24.88 | 15.53 |
| GTE-Qwen2-7B | 36.78 | 51.64 | 51.41 | 38.08 | 17.86 | 28.36 | 23.25 | 15.16 |
| GTE-Base | 34.04 | 48.57 | 50.37 | 37.29 | 16.17 | 26.13 | 19.23 | 12.52 |
| (Using text encoder NV-Embed-v2 at the first stage and GTE-Base at the second stage) | ||||||||
| NVEv2GTE | 37.28 | 51.77 | 48.17 | 37.03 | 18.24 | 28.79 | 24.59 | 15.40 |
| Bold results are better than the NV-Embed-v2 (NVEv2BERT) results above. | ||||||||
Notably, the performance boost in the second-stage CLAP features indicates that the semantic knowledge accumulated in the audio features effectively guides M2D2 to learn useful CLAP features. The second-stage results of LLM-based NV-Embed-v2 and GTE-Qwen2-7B-instruct are 37.24 and 36.78 for R@1 of AudioCaps T-to-A, while the non-LLM-based GTE-Base’s result is 34.04, for example.
The results demonstrate that powerful generalist LLM-based sentence embedding models with SOTA performance in the NLP domain are also useful for learning audio and CLAP features. This finding also suggests that utilizing generalist embedding models in other audio-language tasks would be beneficial, and we leave the possibilities to future work.
The NVEv2GTE results confirm that using a generalist sentence embedding model (GTE-Base) instead of the typically used BERT for the second stage does not improve the final CLAP feature performance. While it performs slightly better than NV-Embed-v2 (NV-Embed-v2BERT) for T-to-A, it degrades more for the AudioCaps A-to-T.
IV-D3 Training design ablations
Training audio encoder in the second stage
While M2D2 freezes the audio encoder in the second stage, an alternative would be to pre-train the audio and text encoder together in this stage, as in [10]. We performed an ablation experiment in which we pre-trained the audio encoder in the second stage. In this case, computational resource constraints decreased the batch size from 2048 to 1024. We also compared the results with M2D2, in which we fine-tuned the audio encoder to AudioSet after the first stage, i.e., the additional stage 1.1 described in Section III-B4.
The results for the audio features shown in Table XII indicate that M2D2, which trains the audio encoder in the second stage, improves the performance of the audio features in the ESC-50 (ESC) and US8K tasks while degrading the performance in other tasks. This trend is the same as M2D2.
On the other hand, Table XIII shows that the audio-text retrieval performance of M2D2 is comparable to M2D2, indicating that training the audio encoder in the second stage does not improve the CLAP feature, which reveals the difference from M2D2. In addition, the decreased batch size may have affected the results.
These experiments suggest that training audio encoders in the second stage, as in the previous study Cacophony[10], might not be a good pre-training strategy for learning representations that perform well for both audio and CLAP features, the goal of M2D2.
| Model | ESC | US8K | SPC | VC1 | VF | CRD | GTZ | NS | SUR | Avg. |
| M2D2, M2D2- | 95.2 | 89.5 | 95.6 | 68.6 | 98.1 | 73.4 | 86.6 | 76.5 | 42.2 | 80.6 |
| M2D2, M2D2 | 97.9 | 89.7 | 94.8 | 56.3 | 97.1 | 69.7 | 86.3 | 76.7 | 41.1 | 78.8 |
| (Section IV-D3: Jointly training text/audio encoders in the second stage) | ||||||||||
| M2D2 | 97.0 | 90.1 | 93.3 | 60.4 | 96.7 | 71.1 | 86.3 | 76.1 | 42.0 | 79.0 |
| (Section IV-D3: Text projector only) | ||||||||||
| M2D2-(TP only) | 94.9 | 88.0 | 94.0 | 60.5 | 97.5 | 71.6 | 83.7 | 76.0 | 41.9 | 78.7 |
| (Section IV-D3: Both audio and text projector, as in previous CLAP models) | ||||||||||
| M2D2-(+TP) | 96.1 | 88.9 | 95.6 | 66.2 | 97.8 | 72.7 | 84.7 | 77.6 | 42.3 | 80.2 |
| AudioCaps | Clotho | |||||||
| Model | T-to-A | A-to-T | T-to-A | A-to-T | ||||
| R@1 | mAP @10 | R@1 | mAP @10 | R@1 | mAP @10 | R@1 | mAP @10 | |
| M2D2 | 37.24 | 52.00 | 52.56 | 38.74 | 18.09 | 28.51 | 24.88 | 15.53 |
| M2D2 | 39.06 | 54.29 | 54.02 | 39.95 | 18.30 | 29.15 | 22.78 | 15.62 |
| (Section IV-D3: Jointly training text/audio encoders in the second stage) | ||||||||
| M2D2 | 37.81 | 52.29 | 50.78 | 37.15 | 17.76 | 28.83 | 21.44 | 14.11 |
| (Section IV-D3: Text projector only) | ||||||||
| M2D2 (TP only) | 34.04 | 48.73 | 45.98 | 34.42 | 17.07 | 27.44 | 22.11 | 14.56 |
| (Section IV-D3: Both audio and text projector, as in previous CLAP models) | ||||||||
| M2D2 (+TP) | 36.07 | 50.52 | 48.69 | 36.93 | 17.53 | 28.09 | 22.58 | 14.72 |
Using text projector only
In this ablation experiment, we removed the audio projector and attached a text projector on top of the text encoder to test the case where an audio feature is learned in the semantic space, and we mapped text features to the space, i.e., an audio feature was learned that works as both conventional and CLAP audio features. We used a one-hidden layer MLP as the text projector.
Adding text projector
This ablation experiment added a text projector to the text encoder, making M2D2 follow a typical CLAP configuration with both audio and text projectors. The results in Tables XII and XIII show that M2D2 (+TP), text projector-added model, degrades performance from M2D2 as in the text projector only ablation. Considering these results, we do not employ a text projector in M2D2.
IV-D4 Visualizations
CLAP audio feature visualizations
We visualize the pre-trained CLAP audio features and show their impact on two downstream tasks with an inherently different nature. We encoded ESC-50 and VoxCeleb1 sounds using M2D2, M2D2 (M2D2 fine-tuned to AudioSet), and CLAP and visualized their features using t-SNE as shown in Fig. 6. We use all 50 class samples from ESC-50 and samples of random ten speakers from VoxCeleb1.
The ESC-50 results show that for all audio features, the samples are clustered by class, indicating that the features are useful for the 50-class classification task. On the other hand, the VoxCeleb1 results show that the CLAP clusters have unclear boundaries compared to M2D2 and M2D2. These observations support the VC1 performance difference in the linear evaluation shown in Table II in Section IV-B1.
The CLAP features are advantageous for ESC-50 since it is an environmental sound classification task, and the pre-training captions contain much information indicating differences in environmental sounds. On the other hand, VoxCeleb1 is a 1251-speaker identification task, and the pre-training captions cannot contain speaker names or IDs, thus making it challenging for conventional CLAP features in the identification task. The conventional CLAP model would learn to represent speaker attributes such as adult or child described in captions, which may explain the tendency not to cluster by individuals. In contrast, M2D2 features form more distinct clusters, indicating that SSL helps learn more information about sound itself, which leads to clearer boundaries for individuals.
These observations demonstrate that jointly learning SSL and CLAP allows M2D2 to include diverse information in CLAP audio features.
Audio projector attention visualizations
Figure 7 shows the attention in the audio projector for the audio samples from AudioCaps and Clotho. The attention map directly shows which patch features the final CLAP audio feature attends to because the audio projector is a single-layer transformer encoder with a single head. We used the M2D2 model and picked the sounds with typical attention maps.
The first example from AudioCaps shows that the projector attends to the louder part of the revving sound and, notably, to the sound characterizing the powering down. The second shows that while the lowest frequency patches attend to speech and music, the patches at the pitch of the whistling attend to each whistling note. The third also shows that the patches at the pitch of the synth sound attend stronger to the corresponding bending up and down or fading in and out. The last example follows the same trend: the lowest frequency patches attend to the sound of waves, while the patches at the seagull calls attend to these sounds. Overall, the observations exemplify that the projector summarizes CLAP audio features based on the patch features that characterize the audio clips.
V Conclusion
This study pursued a general-purpose audio-language representation that provides both general-purpose audio and CLAP features to serve diverse audio applications. For this purpose, we proposed M2D2, a new method combining an SSL M2D and CLAP in a two-stage pre-training and utilizing a recent LLM-based sentence encoder. In the first stage, M2D2 learns a generalizable audio feature using the rich semantic features of the LLM-based sentence encoder, and in the second stage, it learns a CLAP feature using the learned audio feature. While M2D learns highly generalized audio features, the powerful LLM-based embeddings help learn a better CLAP feature through the audio feature. Experiments validated the effectiveness of M2D2 as a general-purpose audio-language representation, highlighted by the SOTA performance of audio features for AudioSet fine-tuning and music tasks and top performance of CLAP features in zero-shot classification, audio-text retrieval, and audio captioning. We conducted ablation studies to analyze M2D2 pre-training and found that two-stage training strategy design and managing the effect of the CLAP objective are vital for learning effective features. We have made our models available online and hope they help the progress of diverse audio studies101010To reviewers, we will update our repository after the notification..

### v2

@@
-M2D2: Exploring General-purpose Audio-Language Representations Beyond CLAP
+M2D-CLAP: Exploring General-purpose Audio-Language Representations Beyond CLAP
 Abstract
-Contrastive language-audio pre-training (CLAP) has addressed audio-language tasks such as audio-text retrieval by aligning audio and text in a common feature space.
-While CLAP addresses general audio-language tasks, its audio features do not generalize well in audio tasks.
-In contrast, self-supervised learning (SSL) models learn general-purpose audio features that perform well in diverse audio tasks.
-We pursue representation learning that can be widely used in audio applications and hypothesize that a method that learns both general audio features and CLAP features should achieve our goal, which we call a general-purpose audio-language representation.
-To implement our hypothesis, we propose M2D2, a second-generation masked modeling duo (M2D) that combines an SSL M2D and CLAP.
-M2D2 learns two types of features using two modalities (audio and text) in a two-stage training process.
-It also utilizes advanced LLM-based sentence embeddings in CLAP training for powerful semantic supervision.
-In the first stage, M2D2 learns generalizable audio features from M2D and CLAP, where CLAP aligns the features with the fine LLM-based semantic embeddings.
-In the second stage, it learns CLAP features using the audio features learned from the LLM-based embeddings.
-Through these pre-training stages, M2D2 should enhance generalizability and performance in its audio and CLAP features.
-Experiments validated that M2D2 achieves effective general-purpose audio-language representation, highlighted with SOTA fine-tuning mAP of 49.0 for AudioSet, SOTA performance in music tasks, and top-level performance in audio-language tasks.
+Contrastive language-audio pre-training (CLAP), which learns audio-language representations by aligning audio and text in a common feature space, has become popular for solving audio tasks.
+However, CLAP’s audio features lack generalizability, whereas self-supervised learning (SSL) models offer general-purpose features that perform well across diverse audio tasks.
+We aim to develop a broadly applicable audio representation and hypothesize that a model that learns both general audio and CLAP features should achieve our goal, which we call a general-purpose audio-language representation.
+To implement our hypothesis, we propose M2D-CLAP, the first approach to jointly learn effective general audio and CLAP features.
+It extends an SSL masked modeling duo (M2D) by incorporating CLAP and utilizes LLM-based sentence embeddings.
+The training process consists of multiple stages.
+In the first stage, generalizable audio features are pre-trained via a multitask objective combining M2D and CLAP, with CLAP leveraging LLM-based semantic embeddings to distill semantic knowledge into them.
+In the following stages, CLAP features are pre-trained and refined with guidance from the learned audio features.
+Experiments demonstrated that M2D-CLAP learns high-performing general audio features (e.g., AudioSet mAP of 49.0, SOTA results in music tasks) and CLAP features, thereby enabling a general-purpose audio-language representation.
 Index Terms:
@@
 In the audio domain, contrastive language-audio pre-training (CLAP) has extended CLIP to audio[2][3][4][5], followed by various advancements[6][7][8][9][10][11] that have exhibited superior performance for applications such as zero-shot classification and audio captioning[7][12][13].
-CLAP can map audio and captions into features aligned in common feature space; therefore, these features can be effective for various downstream application tasks since many task settings are typically related to language, such as inference of class labels.
+CLAP can map audio and captions into features aligned in a common feature space; therefore, these features can be effective for various downstream application tasks since many task settings are typically related to language, such as inference of class labels.
 While the CLAP models have exhibited versatility in audio-language tasks, their audio features do not generalize well in conventional audio tasks because they typically use audio encoders based on supervised learning models.
@@
 For example, they perform well in ESC-50[16], an environmental sound classification similar to the pre-trained AudioSet[17] but poorly in VoxCeleb1[18], a speaker identification task.
-However, some studies have applied CLAP audio features to a task where SSL audio features may perform well, such as [19][20].
+However, some studies have applied CLAP audio features to tasks where SSL audio features may perform well, such as [19][20].
 We pursue representation learning that can be widely effective in diverse audio applications and hypothesize that a generally useful method would be one that learns both general-purpose audio features and CLAP features; we call it a general-purpose audio-language representation.
@@
 In the natural language processing (NLP) domain, sentence embedding models have recently progressed to exploit an LLM[21].
-LLMs have demonstrated overwhelming ability in language-related tasks[22] and have the potential to map audio captions into better semantic features based on knowledge accumulated in huge parameters learned from massive data.
-To implement a general-purpose audio-language representation and achieve its highest potential, we propose M2D2, which extends a SOTA masked prediction-based SSL, M2D[23], combines it with a CLAP objective, and exploits the powerful LLM-based sentence embeddings, as shown in Fig.
+LLMs have demonstrated overwhelming ability in language tasks[22] and have the potential to map audio captions into better semantic features based on knowledge accumulated in huge parameters learned from massive data.
+To implement a general-purpose audio-language representation and achieve its highest potential, we propose M2D-CLAP, the first approach to simultaneously achieve effective general audio and CLAP features.
+It extends a masked prediction-based SSL, M2D[23], combines it with a CLAP objective, and exploits the strong LLM-based sentence embeddings, as shown in Fig.
 1.
-It uses a two-stage pre-training strategy that suppresses required computational resources, thus enabling multitasking of M2D and CLAP, contrastive learning with larger batch size, and even learning from LLM-based sentence embeddings.
-The two-stage pre-training strategy enables learning the best features while avoiding substantial computational costs.
-In the first stage, M2D2 learns from a multitask of M2D and CLAP and uses LLM-based text features for strong semantic supervision, aiming to learn the best general-purpose audio features that can also be projected to an effective CLAP feature.
-In the second stage, M2D2 pre-trains a non-LLM text encoder to learn effective CLAP text features that align better with the learned audio features.
-We aim to exploit the LLM-based knowledge in the sentence embeddings through the stages while avoiding the cost of huge computation by training either audio or text encoder in a stage.
-The experiments validate the effectiveness of M2D2 in achieving our goal of a general-purpose audio-language representation.
-The learned audio feature exhibits SOTA performance in fine-tuning, and the CLAP features demonstrate top performance in audio-language tasks.
-For a successful pre-training by multitasking M2D and CLAP, we find it essential to balance the loss between M2D and CLAP to control the powerful LLM-based semantic supervision.
-The audio projector capacity, LLM-based text encoder, and two-stage training strategy designs are also crucial, which we demonstrate in the ablation studies in Section IV-D.
+M2D-CLAP adopts a multi-stage pre-training strategy that leverages LLM-based sentence embeddings to learn high-quality features while avoiding substantial computational costs.
+In the first stage, M2D-CLAP performs multitask learning with M2D and CLAP objectives to acquire general-purpose audio features aligned with LLM-based sentence embeddings, aiming to distill LLM knowledge into audio representations.
+These features are then fine-tuned on AudioSet to further enhance their effectiveness, particularly for CLAP.
+In the subsequent stages, M2D-CLAP pre-trains a smaller non-LLM text encoder to learn CLAP text features that align with the learned audio features, aiming to transfer the knowledge into the text features.
+By training either the audio or text encoder at each stage, M2D-CLAP achieves efficient training without incurring the high computational cost.
+The experiments validate the effectiveness of M2D-CLAP in achieving our goal of learning a general-purpose audio-language representation.
+The resulting audio and CLAP features show top performance both as frozen and fine-tuned audio features, as well as in audio-language tasks.
+For a successful pre-training by multitasking M2D and CLAP, we find it essential to balance the loss between M2D and CLAP to appropriately control the strong semantic supervision.
+Additionally, the audio projector capacity, the use of an LLM-based text encoder, and the multi-stage training strategy all prove to be critical design choices, as demonstrated in the ablation studies in Section IV-D.
 Our contributions are:
-A general-purpose audio-language representation that provides CLAP features for tasks where audio-language alignment is beneficial and general-purpose audio features for other conventional audio tasks.
-A new pre-training method, M2D2, that combines M2D and CLAP using LLM-based sentence embeddings and provides SOTA general-purpose audio features and top-performing CLAP features, achieving a general-purpose audio-language representation.
-Empirical findings on combining SSL and CLAP in ablation studies and publicly available code with pre-trained weights111https://github.com/nttcslab/m2d for future studies.
-The first study to show that an LLM-based sentence embedding model from the NLP domain helps learn effective audio and CLAP features.
+Introducing a general-purpose audio-language representation that provides general audio and CLAP features to serve a wide range of audio and audio-language tasks.
+Proposing M2D-CLAP, the first (to our knowledge) method to jointly learn top-performing general-purpose audio and CLAP features, thereby realizing a general-purpose audio-language representation.
+Providing empirical findings on combining SSL and CLAP, along with publicly available code and pre-trained weights111https://github.com/nttcslab/m2d to support future research.
+The differences from our conference paper[11] are summarized in the next section.
 II Related Work
 II-A Relationship with our previous work
-This study redefines and extends our previous work, M2D-CLAP[11].
-M2D2 differs from M2D-CLAP in improved audio and CLAP features achieved by the newly employed two-stage pre-training strategy, improved audio projector based on a transformer encoder, and the use of LLM-based sentence embeddings (Section III-B).
-To explore its effectiveness further, we also conducted experiments on audio-text retrieval (Section IV-C2), audio captioning (Section IV-C3), and ablation studies (Section IV-D) in addition to audio transfer learning (Section IV-B) and zero-shot classification (Section IV-C1).
-To make the difference from the final M2D2 clear, we specifically denote M2D-CLAP as M2D20- in the experiments because it is a variant of M2D2-, the first-stage pre-trained M2D2.
+We introduced M2D-CLAP in our previous work[11].
+While it demonstrated the effectiveness, the approach left room for improvement, particularly in the performance of the CLAP features.
+This paper redefines and extends M2D-CLAP to enhance the effectiveness of its audio and CLAP features, leveraging a newly introduced multi-stage pre-training strategy, a transformer-based audio projector, and LLM-based sentence embeddings (Section III-B).
+To further assess its effectiveness, we also conducted experiments on audio-text retrieval, audio captioning, and ablation analyses in addition to transfer learning and zero-shot classification (Sections IV-B)–IV-D).
+To make the difference from the redefined M2D-CLAP clear, we specifically denote it as M2D-CLAP and the previous M2D-CLAP[11] as M2D-CLAP. Note that M2D-CLAP is a variant of the first-stage M2D-CLAP.
 II-B Conventional audio representation models
 Since the release of the large-scale dataset AudioSet[17], various supervised learning models such as PANNs[24], AST[25], and HTS-AT[26] that utilize class labels have followed.
-In contrast, SSL methods such as COLA[27] and BYOL-A[15] are general-purpose audio signal representations that do not depend on task supervision, such as labels.
+In contrast, SSL methods such as COLA[27] and BYOL-A[15] are general-purpose audio representations that do not depend on task supervision, such as labels.
 Masked prediction-based SSLs have progressed and demonstrated remarkable general-purpose performance.
@@
 While they can provide an audio feature as a part of CLAP features, they do not generalize well on conventional audio tasks, as reported in [11][15].
-Methods that take a similar approach to M2D2 are SLIP[36] in the image domain (combines SSL and CLIP), FLAP[8] (combines MAE[37] and CLAP), SupMAM-CLAP[38] (combines Audio-MAE and distillation of a CLAP model), Cacophony[10] (combines Audio-MAE, CLAP, and a captioning model in a two-stage training), and CoLLAT[9] (CLAP with frozen text encoders).
-Unlike these approaches, we combine M2D, CLAP, and the recent LLM-based sentence embeddings in a two-stage pre-training to learn a general-purpose audio-language representation that provides generalizable audio and CLAP features.
+Methods that take a similar approach to M2D-CLAP are SLIP[36] in the image domain (combines SSL and CLIP), FLAP[8] (combines MAE[37] and CLAP), SupMAM-CLAP[38] (combines Audio-MAE and distills a CLAP model), Cacophony[10] (combines Audio-MAE, CLAP, and a captioning model in a two-stage training), and CoLLAT[9] (CLAP with frozen text encoders).
+Unlike these approaches, we combine M2D, CLAP, and the recent LLM-based sentence embeddings in a multi-stage pre-training to learn a general-purpose audio-language representation that provides effective audio and CLAP features.
 III Proposed Method
-We propose M2D2, which extends a masked prediction-based SSL method, M2D, to combine it with CLAP to learn general-purpose audio-language representations effective for a broad range of conventional audio and audio-language tasks.
+We propose M2D-CLAP, which extends M2D to combine it with CLAP to learn general-purpose audio-language representations that provide useful features for a broad range of conventional audio and audio-language tasks.
 III-A Background: Masked Modeling Duo
@@
 III-A2 Transfer learning
-After the pre-training, only the audio encoder is transferred from the online network in downstream tasks.
+After the pre-training, only the audio encoder is transferred from the online network to the downstream tasks.
 As the encoder accepts the input no longer than the predefined duration (e.g., 608 time steps) in the pre-training stage, the inference procedure handles variable-length audio inputs in two ways: one is by cutting the input with the predefined duration and repeatedly encoding the chunks; the other is by interpolating the positional encoding to the input audio duration (i.e., stretching the acceptable input duration).
-The latter is used for fine-tuning the encoder to a task with an input duration similar to the pre-training and the former is used in all other cases, such as linear evaluation.
+The latter is used for fine-tuning the encoder to a task with an input duration similar to the pre-training, and the former is used in all other cases, such as linear evaluation.
 According to the task requirements, the encoder outputs a feature for each patch , which is summarized into a time frame-level feature or an audio clip-level feature .
@@
 As a result, the learned representation reflects both objectives from M2D and the added task.
-III-B Proposed M2D2: M2D Second Generation
-M2D2 is a second-generation M2D, an M2D-X variant that adds a CLAP training objective to implement a general-purpose audio-language representation.
+III-B Proposed M2D-CLAP: M2D Meets CLAP
+M2D-CLAP is an M2D-X variant that combines M2D and CLAP training objectives to implement a general-purpose audio-language representation.
 As shown in Fig.
@@
 The CLAP extension has a semantic network branch that conducts contrastive learning among audio and text features to align them in a semantic feature space.
-M2D2 employs a two-stage training strategy to learn the best of audio and CLAP features under a multitask of M2D and CLAP without the cost of huge computational resources.
+M2D-CLAP employs a multi-stage training strategy to learn the best of audio and CLAP features under a multitask of M2D and CLAP without the cost of huge computational resources.
 In the first stage, it learns effective audio features from M2D that also jointly learns from powerful LLM-based sentence embeddings by aligning the audio features with them through CLAP, aiming to achieve the best audio features that are also effective as CLAP audio features.
 In the second stage, it learns effective CLAP text features that align with the learned audio features in a semantic feature space.
-Through these stages, M2D2 learns effective audio and CLAP features.
+In each stage, it additionally fine-tunes features to achieve effective CLAP performance.
+Through these stages, M2D-CLAP learns effective audio and CLAP features.
 III-B1 First stage pre-training
-In the first stage, the captions are encoded into embeddings in advance using the LLM-based text encoder, and the training process takes audio for M2D input while taking text embeddings for the corresponding caption.
+In the first stage, captions are encoded in advance into embeddings using an LLM-based sentence embedding model, ensuring that the training process receives only audio inputs alongside the corresponding fixed text embeddings.
 M2D then conducts its training process and feeds the audio encoder’s output feature to the semantic network.
-While the original M2D-X framework provides in addition to , M2D2 uses only for training stability.
+While the original M2D-X framework provides in addition to , M2D-CLAP uses only for training stability.
 The audio projector converts the audio features into semantic features .
@@
 We use a powerful LLM-based sentence embedding model and distill its semantic knowledge into the audio encoder through CLAP.
-The audio projector is a transformer encoder to summarize so that the can better represent the semantics of an audio input.
+The audio projector is a transformer encoder to summarize so that the can better represent the semantics of an input audio.
 The training process calculates the CLAP loss using the cosine similarity between and :
@@
 where the loss weights and control the contribution.
+The first-stage pre-training combines supervised learning (CLAP) and SSL (M2D), enabling the audio encoder to learn effective features through their synergy.
+CLAP directs feature learning toward the semantics of the caption text but may miss aspects of sound not explicitly described.
+In contrast, SSL promotes general-purpose feature learning but lacks the guidance provided by textual supervision.
+By integrating both, the model learns semantically aligned yet generalizable representations.
+The learned audio features improve performance on tasks such as environmental sound and music compared to M2D alone, at the cost of slightly reduced performance on speech tasks, but with an overall performance gain.
+The experiments in Section IV-B demonstrate this trade-off.
+Additional stage 1.1: AudioSet fine-tuning
+This stage fine-tunes the first-stage pre-trained audio encoder to AudioSet under a supervised setting using labels.
+This additional process corresponds to adopting the AudioSet fine-tuned audio encoders in the conventional CLAP models[4][5][6].
+Experiments in Sections IV-C2 and IV-C3 show that this fine-tuning step is advantageous with AudioCaps[41] (a subset of AudioSet).
 III-B2 Second stage pre-training
@@
 3.
-To facilitate the CLAP training using a precise loss, we make the audio feature for an audio input consistent by using a low masking ratio and inputting the entire audio sample as is.
+To facilitate CLAP training with a precise loss, we input the entire audio sample without truncation and apply a low masking ratio, allowing a nearly consistent loss calculation.
 In this stage, we employ a modest capacity model, such as BERT, as the text encoder to maintain effective contrastive learning with a sufficient batch size.
@@
 The text encoder encodes captions into the semantic features , and the clap loss is calculated between and , which is the loss for the second stage.
-III-B3 M2D2 transfer learning
+Additional stage 2.1: WavCaps fine-tuning
+This stage further fine-tunes the second-stage audio projector and text encoder to refine the CLAP features using the same dataset as in WavCaps[7] and with no masking of the audio input to learn the best text/audio alignment.
+We show that this process improves performance in audio-language tasks in Section IV-C.
+III-B3 M2D-CLAP transfer learning
 Once pre-trained, we transfer the audio encoder, audio projector, and text encoder as models constituting a general-purpose audio-language representation to the downstream tasks.
@@
 While the audio feature is useful for conventional audio tasks, and work as CLAP features in audio-language tasks such as zero-shot classification and audio-text retrieval.
-III-B4 Additional pre-training stages
-While the two training stages enable the learning of general-purpose features, the experiments show that these features still leave a gap with top performance in existing benchmarks if they are not fine-tuned.
-Therefore, to facilitate the comparison with previous studies, we also use two additional stages to bring the experimental conditions closer to those of existing studies.
-Stage 1.1: AudioSet fine-tuning
-This stage fine-tunes the first-stage pre-trained audio encoder to AudioSet under a supervised setting using labels.
-This additional process corresponds to adopting the AudioSet fine-tuned audio encoders in the conventional CLAP models[4][5][6].
-Experiments in Sections IV-C2 and IV-C3 show that this fine-tuning step is advantageous with AudioCaps[41] (a subset of AudioSet).
-Stage 2.1: WavCaps pre-training
-This stage further pre-trains the second stage audio projector and text encoder to refine the CLAP features using the same dataset as in WavCaps[7] and with no masking of the audio input to learn the best text/audio alignment.
-We show that this process improves performance in audio-language tasks in Section IV-C.
-III-B5 Training design and efficiency
-Though other training strategies are possible, we pre-train the audio encoder first because we aim to exploit the knowledge from the LLM for both audio and text features.
-The M2D2 first stage enables that, unlike typical CLAP models that use a pre-trained audio encoder solely learned from audio data.
-To practically train a CLAP, we need pre-trained audio/text encoders, which is why we train the text encoder in the second stage.
-To enable effective training without a huge computing budget, we train encoders separately and use the LLM offline.
+III-B4 Training design and efficiency
+While other strategies are possible, we pre-train the audio encoder first to leverage LLM knowledge for both audio and text features.
+The first stage of M2D-CLAP enables this, unlike typical CLAP models that use an existing audio encoder pre-trained only on audio data.
+To train a CLAP model in practice, pre-trained audio and text encoders are required; thus, we train the audio encoder first and the text encoder in the second stage.
+To keep training efficient without high computational cost, we train the encoders separately and use the LLM offline.
 We utilize an audio projector only, unlike previous CLAP, which typically uses both audio and text projectors.
-An audio projector is crucial to bridge the gap between the descriptive general audio and abstract (semantic) CLAP features so that they can be learned in a single model.
-A text projector degrades performance in M2D2, as verified in Section IV-D3.
-The staged pre-training allows for training under fewer computing resources, in addition to learning a well-performing representation.
-M2D2 learns from LLM features in the first stage, while the features are constant for the same captions.
-Therefore, M2D2 can substantially reduce resource requirements by pre-computing and retaining features for all captions and referring to pre-computed features instead of loading an LLM instance during training.
+While an audio projector is crucial to bridge the gap between the descriptive general audio and abstract (semantic) CLAP features, a text projector degrades performance in M2D-CLAP, as verified in Section IV-D3.
+The staged pre-training not only enables learning a well-performing representation but also allows training with reduced computational resources.
+In the first stage, M2D-CLAP learns from LLM features that remain fixed for identical captions, enabling a substantial reduction in computational cost by pre-computing features for all captions.
+During training, these pre-computed features are used instead of loading the LLM instance, saving computational resources.
 In contrastive learning, a large batch size is critical for sufficient training.
-While FLAP, a similar method, requires GPUs to achieve a batch size of 2304, M2D2 can conduct training with 4 A100 GPUs (80GB) for a batch size of 2048.
+While FLAP, a similar method, requires GPUs to achieve a batch size of 2304, M2D-CLAP can conduct training with 4 A100 GPUs for a batch size of 2048.
 IV Experiments
-To assess the effectiveness of M2D2 representations, we evaluated the audio features on conventional audio tasks under settings of frozen audio features (Sections IV-B1 and IV-B2) and fine-tuning audio features (Section IV-B3) and the CLAP features using three tasks: zero-shot classification (Section IV-C1), audio-text retrieval (Section IV-C2), and audio captioning (Section IV-C3).
-We also studied how the pre-training design and parameters impact the final performance (Section IV-D), compared the CLAP features using visualizations (Section IV-D4), and show how M2D2 focuses on the audio patches to summarize CLAP audio features (Section IV-D4).
-In the following sections, we refer to the M2D2 models as follows:
-M2D2-: First stage pre-trained representation. Combination of pre-trained audio encoder/audio projector and off-the-shelf LLM-based sentence embedding model as a text encoder.
-M2D2: Second stage pre-trained representation. Combination of the first-stage audio encoder and the second-stage pre-trained audio projector/text encoder.
-M2D2+: Second stage plus stage 2.1 pre-trained representation. Combination of the first-stage audio encoder and the stage 2.1 pre-trained audio projector/text encoder.
-M2D2: Fine-tuned representation to the benchmarks. Combination of the audio encoder fine-tuned in stage 1.1 and the stage 2.1 pre-trained audio projector/text encoder.
-M2D20-: The M2D-CLAP[11] model as an ablation variant of M2D2- for using a non-LLM text encoder and AudioSet only for pre-training audio data.
+To assess the effectiveness of M2D-CLAP representations, we evaluated the audio features on conventional audio tasks under settings of frozen audio features (Sections IV-B1 and IV-B2) and fine-tuning audio features (Section IV-B3) and the CLAP features using three tasks: zero-shot classification (Section IV-C1), audio-text retrieval (Section IV-C2), and audio captioning (Section IV-C3).
+We also studied how the pre-training design and parameters impact the final performance (Section IV-D), compared the CLAP features using visualizations (Section IV-D4), and show how M2D-CLAP focuses on the audio patches to summarize CLAP audio features (Section IV-D4).
+The following sections refer to M2D-CLAP and its intermediate-stage variants used in the ablation studies as follows:
+M2D-CLAP: Final M2D-CLAP representation pre-trained in all two plus two additional stages.
+M2D-CLAP: First stage pre-trained representation. Combination of pre-trained audio encoder/audio projector and off-the-shelf LLM-based sentence embedding model as a text encoder.
+M2D-CLAP: Second-stage pre-trained representation, combining the first-stage audio encoder with the second-stage audio projector/text encoder.
+The audio encoder is not fine-tuned on AudioSet; stage 1.1 is omitted.
+M2D-CLAP: Second stage plus stage 2.1 pre-trained representation, combining the first-stage audio encoder and the stage 2.1 pre-trained audio projector/text encoder.
+M2D-CLAP: The model from the initial study[11] as an ablation variant of M2D-CLAP for using a non-LLM text encoder and AudioSet only for audio data.
 IV-A Experimental setup
 IV-A1 Basic setup
-We followed M2D for the basic setup and implemented M2D2 on top of the M2D codebase as a variant of M2D-X[23].
+We followed M2D for the basic setup and implemented M2D-CLAP on top of the M2D codebase as a variant of M2D-X[23].
 Specifically, the audio encoder is ViT-base[39] with a 768-d feature, input audio duration of 6 s, and patch size of .
@@
 The audio samples were preprocessed to a log-scaled mel spectrogram with a sampling frequency of 16,000 Hz, window size of 25 ms, hop size of 10 ms, and 80 mel-spaced frequency bins in the range of 50 to 8000 Hz.
-We standardized spectrograms with the statistics of the first stage pre-training dataset (e.g., average of and standard deviation of ).
+We standardized spectrograms with the statistics of the first stage dataset (e.g., average of and standard deviation of ).
 IV-A2 Pre-training setup
@@
 In the additional stage 1.1, we fine-tuned the first-stage audio encoder in the same way as in the fine-tuning experiments in Section IV-B3.
-In the additional stage 2.1, we pre-trained the second-stage M2D2 with the same setting as in the second stage once again, except for a masking ratio of 0.0 (i.e., no masking) and a limited dataset.
+In the additional stage 2.1, we used the same setting as in the second stage once again, except for a masking ratio of 0.0 (i.e., no masking) and a limited dataset.
 IV-A3 Pre-training datasets
@@
 Since the curated dataset included data samples for the downstream tasks (AudioCaps[41], Clotho[47], ESC-50[16], US8K[48]) used in the evaluations, we removed all of these data samples.
-For the audio data, we randomly cropped 6-s segments during the training.
+For training audio, we randomly cropped 6-s segments.
 The second stage
@@
 | †FT, LE, and ZS stand for fine-tuning, linear evaluation, and zero-shot classification, respectively. | ||||||||||||
-under a unified linear evaluation except Pengi, Cacophony, and CoLLAT.
 | Audio enc. fine-tuning on AudioSet | Env. sound tasks | Speech tasks | Music tasks | ||||||||
@@
 | ATST-Frame [32] | 90.9 0.6 | 85.8 | 94.9 | 77.4 | 98.8 0.3 | 72.3 0.7 | 82.9 6.0 | 75.9 | 40.6 0.2 | 79.9 | |
+| MATPAC [56] | 93.5 | 89.4 | - | - | - | - | 85.9 | 74.6 | - | - | |
 | Cacophony (stage 1)[10] | 87.0 | - | 92.2 | - | - | 69.7 | 83.8 | - | - | - | |
@@
 | PANNs CNN14[24] | ✓ | 90.70.8 | 81.90.1 | 51.80.3 | 7.80.1 | 75.10.2 | 51.81.3 | 78.73.0 | 65.90.2 | 10.90.3 | 57.2 |
-| (Previous studies: CLAP and audio-language models) | |||||||||||
+| (Previous studies: Audio-language models) | |||||||||||
 | LAION-CLAP [4] | ✓ | 97.3 0.5 | 86.9 0.5 | 75.9 0.5 | 13.4 0.4 | 80.3 0.2 | 54.6 1.0 | 84.3 2.6 | 72.2 1.1 | 14.8 0.5 | 64.4 |
@@
 | M2D[23] | ✓ | 96.7 0.5 | 89.3 0.1 | 93.2 0.2 | 52.7 0.6 | 96.9 0.0 | 69.5 1.8 | 87.6 0.9 | 75.6 0.6 | 40.5 0.4 | 78.0 |
-| (Proposed: Audio-language models) | |||||||||||
-| M2D2, M2D2-, M2D2+ | 95.2 0.3 | 89.5 0.2 | 95.6 0.1 | 68.6 0.5 | 98.1 0.1 | 73.4 0.7 | 86.6 1.6 | 76.5 0.4 | 42.2 0.3 | 80.6 | |
-| M2D2 | ✓ | 97.9 0.5 | 89.7 0.3 | 94.8 0.2 | 56.3 0.5 | 97.1 0.0 | 69.7 0.3 | 86.3 1.8 | 76.7 0.4 | 41.1 0.3 | 78.8 |
-| (Ablation using a non-LLM text encoder and AudioSet only as audio data. The underlined results outperform M2D2-) | |||||||||||
-| M2D20-(M2D-CLAP[11]) | 96.3 0.3 | 88.8 0.6 | 95.8 0.3 | 70.3 0.4 | 98.3 0.1 | 73.4 0.2 | 84.1 1.5 | 78.0 0.5 | 42.4 0.6 | 80.8 | |
+| (Proposed: Audio-language model) | |||||||||||
+| M2D-CLAP | ✓ | 97.9 0.5 | 89.7 0.3 | 94.8 0.2 | 56.3 0.5 | 97.1 0.0 | 69.7 0.3 | 86.3 1.8 | 76.7 0.4 | 41.1 0.3 | 78.8 |
+| (Ablations: No fine-tuning to AudioSet. M2D-CLAP is pre-trained using a non-LLM text encoder and audio data only from AudioSet.) | |||||||||||
+| M2D-CLAP | 95.2 0.3 | 89.5 0.2 | 95.6 0.1 | 68.6 0.5 | 98.1 0.1 | 73.4 0.7 | 86.6 1.6 | 76.5 0.4 | 42.2 0.3 | 80.6 | |
+| M2D-CLAP[11] | 96.3 0.3 | 88.8 0.6 | 95.8 0.3 | 70.3 0.4 | 98.3 0.1 | 73.4 0.2 | 84.1 1.5 | 78.0 0.5 | 42.4 0.6 | 80.8 | |
 | Results quoted from corresponding papers when they are better than ours or unavailable in our test. | |||||||||||
@@
 AudioSet has two settings: AS2M with all 2M training samples and AS20K with 21K samples from the balanced train segments, and Surge is a non-semantic pitch audio classification of 88 MIDI notes.
-IV-B Evaluation: Audio Features
-We evaluated the effectiveness of the pre-trained audio features under the frozen feature and fine-tuning settings on general tasks and the frozen feature setting on music tasks.
-We followed [15][23] for the frozen feature (linear evaluation) and fine-tuning experimental setup using a unified evaluation platform, EVAR444https://github.com/nttcslab/eval-audio-repr, on general tasks, and used MARBLE benchmark[59] for the music task evaluation.
-IV-B1 Frozen audio features on general audio tasks
-Experimental setup
-We followed the standard linear evaluation procedure[15][23].
-Specifically, the evaluation pipeline first converts downstream task samples into features using a frozen pre-trained model, trains a linear layer with task labels, and then conducts tests for the trained linear layer.
-We used the validation set for early stopping with patience of 20 epochs and trained the linear layer for up to 200 epochs using the Adam optimizer and a learning rate of 0.00003.
-Results: Frozen audio features on general audio tasks
-The results of the linear evaluation shown in Table II indicate that supervised learning and CLAP models perform poorly on some tasks, whereas the M2D2/M2D variants generally perform well.
-For example, supervised learning models AST, HTS-AT, and PANNs show 7.8 to 18.1 on VC1, and CLAP models such as LAION-CLAP show 8.4 to 27.2 on Surge, whereas the M2D variants show 52.7 to 73.4 on VC1 and 40.5 to 42.4 on Surge; we also observe these differences in other speech tasks.
-Note that the CLAP models use supervised learning models as their audio encoders; hence, the results indicate that applying audio features learned using labels/captions to tasks outside the coverage of these signals is challenging.
-For example, a training signal of ”Male speech, man speaking” does not cover who is speaking on VC1 (speaker identification).
-BEATs and CED, pre-trained with distillation of the previous training cycle models, show a similar trend to supervised learning models.
-CED shows 35.2 on VC1, and BEATs shows 35.7 on Surge, for example.
-These models have distilled the pre-trained models on AudioSet, making them focused more on AudioSet data distribution.
-The results also indicate that their features become difficult to apply for tasks with different data distributions.
-In contrast, M2D2/M2D variants and an SSL model ATST perform well in all tasks.
-M2D2 improves performance on ESC-50 at 94.9 and US8K at 89.1 compared to M2D (note that M2D2-/M2D2+ share the same audio encoder with M2D2).
-However, it exhibits slight performance degradation in the speech task, especially VC1, indicating a trade-off between M2D and CLAP’s training signal.
-Notably, M2D2 (M2D2 encoder fine-tuned to AudioSet) does not degrade performance severely in speech tasks while improving performance on ESC-50 and US8K.
-It shows SOTA performance of 97.9 on ESC-50 and 89.7 on US8K.
-However, it degrades performance to 56.3 on VC1 from M2D2’s 68.6, while it is still in the good performance range compared to previous models.
-Similarly, M2D, a fine-tuned version of the original M2D, improves performance on ESC-50/US8K and degrades on others slightly more than M2D2 but not as severely as the supervised learning models.
-These results show that the M2D pre-training helps learn general-purpose representations that do not severely compromise performance against subsequent fine-tuning.
-In addition, M2D2, which is pre-trained only on AudioSet using GTE-Base[60] (non-LLM-based sentence embedding model) as a text encoder, performs better than M2D2 in most tasks, indicating that the pre-training setting, especially relying solely on AudioSet is beneficial for performing well with frozen features.
-However, as shown in Section IV-C, its CLAP features severely underperform M2D2.
-In summary, the experiments validated the effectiveness of M2D2 and the variants as a frozen audio extractor on general audio tasks.
+All tasks are classification setups, and results are accuracies or mean average precision (mAP).
+When data contained variable-length audio, we randomly cropped or zero-padded the average audio length for each sample to create a batch of length-aligned audio fed into the model as in [15][23].
+We also used AudioCaps[41] and Clotho[47] for audio-text retrieval and audio captioning. In addition, we used various music tasks supported by the MARBLE benchmark[60].
 | MTT | GTZAN | EMO | NSynth | VocalSet | MTG | ||||||||||||
@@
 | Model | ROC | AP | Acc | R2V | R2A | Acc | Acc | Acc | Acc | ROC | AP | ROC | AP | ROC | AP | ROC | AP |
-| (Previous studies: Audio and audio-language models) | |||||||||||||||||
+| (Previous studies: General audio and audio-language models) | |||||||||||||||||
 | ATST-Frame[32] | 91.2 | 39.7 | 79.0 | 60.3 | 76.7 | 77.0 | 84.2 | 76.8 | 88.7 | 78.5 | 21.0 | 77.0 | 15.1 | 87.0 | 19.5 | 83.0 | 29.7 |
@@
 | M2D | 91.5 | 40.9 | 86.1 | 59.0 | 73.0 | 78.7 | 88.2 | 76.8 | 91.3 | 79.7 | 23.3 | 77.8 | 15.6 | 87.5 | 20.3 | 83.4 | 30.6 |
-| (Proposed: Audio-language models) | |||||||||||||||||
-| M2D2, M2D2-, M2D2+ | 91.5 | 40.8 | 82.6 | 59.3 | 76.7 | 77.4 | 89.1 | 77.4 | 91.8 | 79.3 | 21.7 | 78.0 | 15.9 | 87.9 | 20.8 | 83.7 | 30.3 |
-| M2D2 | 91.8 | 41.6 | 85.9 | 58.2 | 76.2 | 79.7 | 89.5 | 78.9 | 92.7 | 79.7 | 22.2 | 78.8 | 15.8 | 88.3 | 20.9 | 84.0 | 30.9 |
-| (Ablation using a non-LLM text encoder and AudioSet only as audio data. The underlined results outperform M2D2-) | |||||||||||||||||
-| M2D20-(M2D-CLAP[11]) | 91.6 | 41.0 | 84.3 | 61.9 | 77.4 | 80.6 | 88.8 | 78.1 | 92.3 | 79.0 | 21.6 | 78.2 | 16.0 | 88.0 | 20.8 | 83.8 | 30.4 |
+| (Proposed: General audio-language model) | |||||||||||||||||
+| M2D-CLAP | 91.8 | 41.6 | 85.9 | 58.2 | 76.2 | 79.7 | 89.5 | 78.9 | 92.7 | 79.7 | 22.2 | 78.8 | 15.8 | 88.3 | 20.9 | 84.0 | 30.9 |
+| (Ablations: No fine-tuning to AudioSet. M2D-CLAP is pre-trained using a non-LLM text encoder and audio data only from AudioSet.) | |||||||||||||||||
+| M2D-CLAP | 91.5 | 40.8 | 82.6 | 59.3 | 76.7 | 77.4 | 89.1 | 77.4 | 91.8 | 79.3 | 21.7 | 78.0 | 15.9 | 87.9 | 20.8 | 83.7 | 30.3 |
+| M2D-CLAP[11] | 91.6 | 41.0 | 84.3 | 61.9 | 77.4 | 80.6 | 88.8 | 78.1 | 92.3 | 79.0 | 21.6 | 78.2 | 16.0 | 88.0 | 20.8 | 83.8 | 30.4 |
+IV-B Evaluation: Audio Features
+We evaluated the effectiveness of the pre-trained audio features under the frozen feature and fine-tuning settings on general tasks and the frozen feature setting on music tasks.
+We followed [15][23] for the frozen feature (linear evaluation) and fine-tuning experimental setup using a unified evaluation platform, EVAR444https://github.com/nttcslab/eval-audio-repr, for general tasks.
+We also evaluated features for music tasks using the MARBLE benchmark[60].
+IV-B1 Frozen audio features on general audio tasks
+Experimental setup
+We followed the standard linear evaluation procedure[15][23].
+Specifically, the evaluation pipeline first converts downstream task samples into features using a frozen pre-trained model, trains a linear layer with task labels, and then conducts tests for the trained linear layer.
+We used the validation set for early stopping with patience of 20 epochs and trained the linear layer for up to 200 epochs using the Adam optimizer and a learning rate of 0.00003.
+Results: Frozen audio features on general audio tasks
+The linear evaluation results in Table II show that M2D and M2D-CLAP variants perform well across the tasks, whereas supervised learning models (AST, HTS-AT, and PANNs) and CLAP models (e.g., LAION-CLAP) perform poorly on speech tasks and Surge.
+For instance, their scores range from 7.8 to 18.1 on VC1 for supervised models and 8.4 to 27.2 on Surge for CLAP models.
+In contrast, M2D and M2D-CLAP variants achieve significantly higher scores, ranging from 52.7 to 73.4 on VC1 and from 40.5 to 42.4 on Surge.
+Note that the CLAP models employ backbone audio encoders trained with supervised learning, which limits the generalizability of their audio features to tasks outside the scope of their training labels and captions.
+For example, a label such as ”Male speech, man speaking” does not capture speaker identity, as required in VC1 (speaker identification task).
+BEATs and CED, pre-trained with distillation of the previous training cycle models, show a similar trend to supervised learning models.
+CED shows 35.2 on VC1, and BEATs shows 35.7 on Surge, for example.
+These models have distilled the pre-trained models on AudioSet, making them focused more on AudioSet data distribution.
+This specialization may hinder their generalization to tasks involving different data distributions.
+In contrast, M2D, M2D-CLAP, and an SSL model, ATST, perform well in all tasks, and M2D-CLAP achieves SOTA performance on ESC-50 at 97.9 and US8K at 89.7.
+However, fine-tuning on AudioSet may have reduced the sensitivity of the representations to fine-grained speaker differences, causing a performance drop in M2D-CLAP on VC1 (56.3 compared to 73.4 with M2D), although 56.3 still remains a competitive performance.
+The first-stage model, M2D-CLAP, also shows slightly lower performance on speech tasks and reduces the overall average performance compared to M2D, suggesting that the use of AudioSet captions or label supervision may be the cause.
+In addition, M2D-CLAP, which is pre-trained only on AudioSet using GTE-Base[64] (non-LLM-based sentence embedding model) as a text encoder, performs better than M2D-CLAP in most tasks, indicating that the pre-training setting, especially relying solely on AudioSet is beneficial for performing well with frozen audio features.
+However, as shown in Section IV-C, its CLAP features severely underperform M2D-CLAP.
+In summary, M2D-CLAP maintains strong performance as a frozen audio feature extractor on general audio tasks, albeit with compromised performance on speech tasks.
 IV-B2 Frozen audio features on music tasks
-To further assess the frozen feature, we evaluated models on the music audio tasks with various music audio signals.
+To further assess the frozen feature, we evaluated models on the music audio tasks with various types of audio signals.
 Experimental setup
-We used the MARBLE benchmark[59], which features diverse music audio tasks for music tracks, instrument sounds, and vocals.
-We used the tasks for which we could reproduce the results in the paper with the released code555https://github.com/a43992899/MARBLE-Benchmark, including music tagging and genre classification on MagnaTagATune (MTT)[64], GTZAN[53], and MTG-Jamendo (MTG)[65], emotion detection on Emomusic[66] and MTG, pitch and instrument (inst) classification on NSynth[54] and MTG, and vocal technique detection (tech) and singer identification on VocalSet[67].
+We used the MARBLE benchmark[60], which covers diverse music audio tasks for music tracks, instrument sounds, and vocals.
+We selected the tasks for which we could reproduce the results with the released code555https://github.com/a43992899/MARBLE-Benchmark.
+The tasks include music tagging and genre classification on MagnaTagATune (MTT)[65], GTZAN[53], and MTG-Jamendo (MTG)[66], emotion detection on Emomusic[67] and MTG, pitch and instrument (inst) classification on NSynth[54] and MTG, and vocal technique detection (tech) and singer identification on VocalSet[68].
 The evaluation metrics are the area under the ROC curve (ROC) and the average precision (AP), except for Emomusic, which uses the determination coefficient between the model’s regression results and human annotations of valence (R2V) and arousal (R2A)[67].
@@
 These facts would explain the comparable performance of general models on the music-specific benchmark.
-However, the CLAP model, CLAP2023, using a supervised learning audio backbone, underperforms other general and music models except for NSynth Inst666NSynth is contained in the CLAP2023 training data., suggesting the need for SSL to generalize in music tasks.
-In contrast, the SSL general audio models perform well in music tasks, indicating that pre-training on AudioSet with SOTA SSL methods can generalize well.
-The models pre-trained using both SSL and supervised learning of AudioSet (M2D, M2D2, and BEATs) show stronger performance, and M2D2 demonstrates new SOTA results over the SOTA music audio models.
-In addition, as in the frozen audio feature evaluation on general tasks, M2D2 performs better than M2D2 in most music tasks, also indicating that pre-training solely on AudioSet is beneficial for performing well with frozen features.
-To summarize, the experiment confirmed that M2D2 models perform well and show SOTA performance in diverse music tasks.
+However, CLAP2023, which uses a supervised learning-based audio backbone, underperforms other general and music models (except NSynth (Inst)666The CLAP2023 training data contains NSynth.), suggesting that SSL on AudioSet is important for generalizing to music tasks, as demonstrated by other SSL-based general audio models.
+The models pre-trained using both SSL and supervised learning of AudioSet (M2D-CLAP and BEATs) show stronger performance, and M2D-CLAP demonstrates new SOTA results on MTT, VocalSet, and MTG.
+However, these models underperform music audio models on the NSynth (Pitch) task, indicating that music models have higher resolution in capturing frequency-related information than general models.
+Notably, M2D-CLAP achieves performance on par with MULE, a music model, on this task.
+In addition, as in the evaluation on general tasks in Table II, M2D-CLAP performs better than M2D-CLAP in most tasks, also indicating that pre-training solely on AudioSet is beneficial for performing well with frozen features.
+To summarize, the experiment confirmed that M2D-CLAP performs well and shows SOTA performance in music tasks.
 IV-B3 Fine-tuning
+For fine-tuning, we focus primarily on evaluating M2D-CLAP. We discuss the final M2D-CLAP as an ablation model because it has already been fine-tuned to AudioSet once in the additional stage 1.1.
 Experimental setup
@@
 We used the hyperparameters for each task as listed in Table IV.
-For Structured Patchout, we used 768-d features calculated by averaging all ViT output features because Eq.
+For Structured Patchout, we used 768-d features calculated by averaging all ViT outputs because Eq.
 (2) is not applicable to the masked patches.
 We used the tasks commonly used in previous studies, namely ESC-50, SPCV2, and VC1, the same as in the linear evaluation, plus AS2M and AS20K.
+Unlike the linear evaluation covering most models, we limited fine-tuning evaluation to our models, as reproducing previous best results often involves non-trivial tuning.
 | Parameter | AS2M | AS20K | ESC-50 | SPCV2 | VC1 |
@@
 | Optimizer | LARS | SGD | SGD | SGD | AdamW |
-| Mixup ratio | 0.5 | 0.3 | 0.0 | 0.3 | 0.0 |
+| Mixup[70, 15] ratio | 0.5 | 0.3 | 0.0 | 0.3 | 0.0 |
 | Random resize crop | - | ✓ | ✓ | ✓ | - |
@@
 | M2D | 47.9 0.0 | 38.6 0.1 | 96.0 0.2 | 98.4 0.1 | 96.3 0.2 |
-| (Proposed: Audio-Language models) | |||||
-| M2D2, M2D2-, M2D2+ | 49.0 0.1 | 42.1 0.0 | 97.8 0.1 | 98.4 0.0 | 95.5 0.3 |
-| M2D2 | - | 46.7 0.2 | 98.5 0.0 | 98.4 0.1 | 94.0 0.2 |
-| (Ablation using a non-LLM text encoder and AudioSet only as audio data) | |||||
-| M2D20-(M2D-CLAP[11]) | 48.5 0.1 | 41.8 0.2 | 97.4 0.2 | 98.3 0.1 | 95.5 0.2 |
+| (Proposed: Audio-Language model) | |||||
+| M2D-CLAP | 49.0 0.1 | 42.1 0.0 | 97.8 0.1 | 98.4 0.0 | 95.5 0.3 |
+| (Ablations: M2D-CLAP have been fine-tuned to AudioSet in stage 1.1.) | |||||
+| M2D-CLAP | - | 46.7 0.2 | 98.5 0.0 | 98.4 0.1 | 94.0 0.2 |
+| M2D-CLAP[11] | 48.5 0.1 | 41.8 0.2 | 97.4 0.2 | 98.3 0.1 | 95.5 0.2 |
 | (Reference: Models using ensemble of audio models or distillation from large models) | |||||
@@
 Results: Fine-tuning
-The experimental results in Table V show that M2D2 variants further improve the performance of the baseline M2D.
-Unlike the linear evaluation that evaluated all models except a few, we evaluated only our models in fine-tuning because reproducing the best results of the previous methods is quite demanding.
-M2D2 improves the performance from M2D, which shows SOTA results of 49.0 for AS2M and top-level results of 42.1 for AS20K and 97.8 for ESC-50.
-On the other hand, VC1 performance is 95.5, which is lower than the original M2D’s performance of 96.3, suggesting a trade-off for learning CLAP.
-M2D2 with an audio encoder fine-tuned to AudioSet exhibits an even stronger trend, showing SOTA results of 46.7 for AS20K and 98.5 for ESC-50 while 94.0 for VC1, a -1.5 pp degradation compared to M2D2.
-The experiment confirmed that M2D2 achieves audio features that exhibit SOTA performance.
-Notably, it shows top performance with a single pre-training, unlike other SOTA models that require multi-iteration/model pre-training.
-In addition, the performance boost of M2D2 from M2D20- (M2D-CLAP) suggests the benefit of using a powerful LLM-based sentence encoder.
+Table V shows that M2D-CLAP improves upon the baseline M2D, achieving SOTA performance on AS2M (49.0) and strong results on AS20K (42.1) and ESC-50 (97.8), outperforming the previous top model, BEATsiter3+, except on ESC-50.
+However, VC1 performance degrades to 95.5, lower than the original M2D’s performance (96.3), suggesting that learning from the CLAP supervision sacrifices fine-grained feature representation.
+Compared to ablation M2D-CLAP, M2D-CLAP shows superior performance on all tasks, suggesting that the effect of the semantic supervision of LLM-based sentence embeddings is effective in fine-tuning, but does not contribute as a frozen feature, as confirmed in Section IV-B1.
+The final M2D-CLAP, which we handle as ablation, shows improved performance of 46.7 for AS20K and 98.5 for ESC-50, while 94.0 for VC1, a -1.5 pp degradation compared to M2D-CLAP.
+The degradation of VC1 reaffirms the trade-off caused by fine-tuning under AudioSet’s supervision.
+In summary, M2D-CLAP provides top-performing audio features for fine-tuning.
+Notably, this is achieved by a single pre-training process, unlike other SOTA models that require multiple stages or models.
+The first stage model M2D-CLAP also shows consistent performance gains across other experiments in Section IV-B.
 IV-C Evaluation: CLAP Features
@@
 We conducted experiments to evaluate the zero-shot classification performance of CLAP features.
-Under the zero-shot classification setting, a model infers the class label of audio by picking a class with the highest similarity between a CLAP audio feature and CLAP text feature of each class.
+In this setting, the model, without any task-specific training, predicts the class label of an audio clip by selecting the one whose CLAP text feature has the highest similarity to the CLAP audio feature.
 Experimental setup
 We chose the tasks found in previous papers, which include ESC-50 (ESC), US8K, CREMA-D (CRD), GTZAN (GTZ), NSynth (NS), AudioSet (AS), and FSD50K (FSD).
-We used AudioSet and FSD50K as references because M2D2 was pre-trained on AudioSet and part of FSD50K in WavCaps.
+We used AudioSet and FSD50K as references because M2D-CLAP was pre-trained on AudioSet and partially on FSD50K in WavCaps.
 We conducted experiments using a standard procedure to infer a class label of audio with the closest cosine distance between the features of each test sample and the label’s caption, and we obtained accuracies from the inference results.
 Table VI summarizes the rules for converting task labels to the captions.
-We used the evaluation platform EVAR, as in Section IV-B, to evaluate M2D2 and pervious models.
-Results: Zerio-shot-classification
+We used the evaluation platform EVAR, as in Section IV-B.
+Results: Zero-shot classification
 Table VII shows the experimental results.
-We compare results based on a strict zero-shot classification setting as in [7]; we only include experimental results that do not use data from downstream tasks during training in our comparisons and gray out all other results as references.
-Results show that M2D2 variants achieve top-level performance on all tasks.
-Among the variants, M2D2- outperforms previous models with a SOTA result of 79.31 for GTZAN, while the second-stage pre-training degrades performance, such as 72.41 with M2D2.
-M2D2+ brings the ESC-50 and US8K results to the top level, indicating that CLAP training with the limited combination of WavCaps, AudioCaps, and Clotho is advantageous for these tasks compared to M2D2 and M2D2- that use large-scale datasets.
-Furthermore, M2D2- significantly improves performance on these tasks from M2D20-, indicating the benefit of using an LLM-based text encoder.
-On the other hand, M2D2 improves CREMA-D closer to the top, CLAP2023.
-In summary, we validated that M2D2 achieves top-level performance in the zero-shot classification setting.
+We compare results based on a strict zero-shot classification setting as in [7]; we only include experimental results that do not use data from downstream tasks during training in our comparisons, and all other results are grayed out as references.
+Results show that M2D-CLAP variants perform well on all tasks, and M2D-CLAP achieves top-level performance in all tasks.
+Among the ablation variants, M2D-CLAP outperforms previous models with a SOTA result of 79.31 for GTZAN, while the second-stage pre-training degrades performance, such as 72.41 with M2D-CLAP.
+M2D-CLAP brings the ESC-50 and US8K results to the top level, indicating that CLAP training with the limited combination of WavCaps, AudioCaps, and Clotho is advantageous for these tasks compared to the models M2D-CLAP that use large-scale datasets.
+Furthermore, M2D-CLAP significantly improves performance on these tasks from M2D-CLAP[11], indicating the benefit of using an LLM-based text encoder as well as incorporating additional data from VGGSound and WavCaps.
+In summary, we validated that M2D-CLAP achieves top-level performance in the zero-shot classification setting.
 | Task | Rule |
@@
 | CoLLAT[9] | 9 | 19 | 84 | 77 | - | - | - |
+| GLAP[78] | - | 40.9 | 88.8 | 78.9 | - | 69.6 | 31.3 |
 | JMLA [79] | - | - | - | - | - | 64.82 | - |
-| (Proposed models) | |||||||
-| M2D2- | 23.15 | 43.43 | 90.07 | 76.52 | 17.41 | 79.31 | 31.76 |
-| M2D2 | 28.27 | 46.76 | 93.67 | 81.92 | 23.50 | 72.41 | 24.57 |
-| M2D2+ | 27.13 | 42.03 | 94.55 | 82.90 | 21.73 | 70.86 | 29.52 |
-| M2D2 | 30.05 | 50.04 | 94.30 | 82.30 | 28.56 | 70.69 | 30.57 |
-| (Ablation using a non-LLM text encoder and AudioSet only as audio data) | |||||||
-| M2D20-(M2D-CLAP) | 27.24 | 40.82 | 75.45 | 72.40 | 17.73 | 75.17 | 23.39 |
+| (Proposed model) | |||||||
+| M2D-CLAP | 30.05 | 50.04 | 94.30 | 82.30 | 28.56 | 70.69 | 30.57 |
+| M2D-CLAP | 23.15 | 43.43 | 90.07 | 76.52 | 17.41 | 79.31 | 31.76 |
+| M2D-CLAP | 28.27 | 46.76 | 93.67 | 81.92 | 23.50 | 72.41 | 24.57 |
+| M2D-CLAP | 27.13 | 42.03 | 94.55 | 82.90 | 21.73 | 70.86 | 29.52 |
+| M2D-CLAP[11] | 27.24 | 40.82 | 75.45 | 72.40 | 17.73 | 75.17 | 23.39 |
 | (Reference studies that employ additional techniques on top of existing CLAP models.) | |||||||
@@
 | FLAP[8] | LA+A+C | 41.5 | 75.5 | 86.0 | - | 53.0 | 84.1 | 92.6 | - | 20.3 | 46.5 | 58.8 | - | 25.5 | 53.4 | 67.9 | - | |
-| (Proposed models) | ||||||||||||||||||
-| M2D2 | LS+WC | 27.34 | 58.14 | 71.45 | 40.14 | 36.05 | 65.31 | 77.43 | 23.95 | 15.92 | 37.88 | 49.26 | 25.37 | 18.09 | 39.90 | 52.54 | 12.39 | |
-| M2D2 | LS+WC+A+C | 37.24 | 72.52 | 84.54 | 52.00 | 52.56 | 82.34 | 90.60 | 38.74 | 18.09 | 42.53 | 55.52 | 28.51 | 24.88 | 48.13 | 60.48 | 15.53 | |
-| M2D2+ | WC+A+C | 40.54 | 74.52 | 86.92 | 55.09 | 57.99 | 83.91 | 91.43 | 41.72 | 19.43 | 45.84 | 58.64 | 30.55 | 26.32 | 51.20 | 65.07 | 17.65 | |
-| M2D2 | WC+A+C | ✓ | 41.96 | 77.07 | 88.57 | 56.74 | 59.25 | 84.74 | 92.79 | 43.24 | 20.08 | 45.99 | 59.48 | 31.21 | 24.98 | 51.67 | 64.59 | 17.35 |
-| (Ablations using a non-LLM text encoder. For the audio data, M2D20-(M2D-CLAP) used AudioSet only while +WC used AudioSet and WavCaps) | ||||||||||||||||||
-| M2D20-(M2D-CLAP) | LS | 24.97 | 58.58 | 72.50 | 39.00 | 25.50 | 58.93 | 76.07 | 18.59 | 10.24 | 26.49 | 37.07 | 17.33 | 11.00 | 25.93 | 35.31 | 7.20 | |
+| GLAP[78] | LS+WC+A+C | 41.7 | - | 86.1 | - | 54.4 | - | 91.1 | - | 19.4 | - | 58.3 | - | 21.8 | - | 61.5 | - | |
+| (Proposed model) | ||||||||||||||||||
+| M2D-CLAP | WC+A+C | ✓ | 41.96 | 77.07 | 88.57 | 56.74 | 59.25 | 84.74 | 92.79 | 43.24 | 20.08 | 45.99 | 59.48 | 31.21 | 24.98 | 51.67 | 64.59 | 17.35 |
+| M2D-CLAP | LS+WC | 27.34 | 58.14 | 71.45 | 40.14 | 36.05 | 65.31 | 77.43 | 23.95 | 15.92 | 37.88 | 49.26 | 25.37 | 18.09 | 39.90 | 52.54 | 12.39 | |
+| M2D-CLAP | LS+WC+A+C | 37.24 | 72.52 | 84.54 | 52.00 | 52.56 | 82.34 | 90.60 | 38.74 | 18.09 | 42.53 | 55.52 | 28.51 | 24.88 | 48.13 | 60.48 | 15.53 | |
+| M2D-CLAP | WC+A+C | 40.54 | 74.52 | 86.92 | 55.09 | 57.99 | 83.91 | 91.43 | 41.72 | 19.43 | 45.84 | 58.64 | 30.55 | 26.32 | 51.20 | 65.07 | 17.65 | |
+| (Ablations using a non-LLM text encoder. For the training data, M2D-CLAP[11] used AudioSet only while +WC used AudioSet and WavCaps.) | ||||||||||||||||||
+| M2D-CLAP[11] | LS | 24.97 | 58.58 | 72.50 | 39.00 | 25.50 | 58.93 | 76.07 | 18.59 | 10.24 | 26.49 | 37.07 | 17.33 | 11.00 | 25.93 | 35.31 | 7.20 | |
 | +WC | LS+WC | 26.06 | 57.89 | 72.64 | 39.79 | 28.63 | 61.65 | 75.97 | 19.69 | 14.81 | 35.44 | 47.00 | 23.78 | 15.41 | 33.78 | 47.27 | 10.42 | |
 IV-C2 Audio-Text Retrieval
-We evaluated the effectiveness of the CLAP features in the audio-text retrieval task setting.
-This task involves searching for text from audio or audio from text, so unlike zero-shot classification, it requires fine-grained comparisons to find pairs at the data instance level[80].
-Therefore, the more accurately the features represent the audio and text, the better the task performance.
+The audio-text retrieval task evaluates CLAP features and involves searching for text from audio or audio from text, so unlike zero-shot classification, it requires fine-grained comparisons to find pairs at the data instance level[82].
+Therefore, task performance improves as the features more accurately align the audio and text.
 Experimental setup
@@
 Results: Audio-text retrieval
-Table VIII shows that M2D2- improves overall performance compared to M2D20- and M2D20-+WC, which is pre-trained on the same dataset as M2D2- to make the difference only in the text encoder and audio projector used.
-M2D2- improves performance on most metrics, e.g., AudioCaps Text-to-Audio R@1 performance of 27.34 compared to M2D20-+WC’s 26.06.
-The results confirm the effectiveness of the M2D2- updates from M2D20- (M2D-CLAP) using an LLM-based text (sentence) encoder and transformer-based audio projector.
-M2D2 further improves overall performance compared to M2D2-, by pre-training the text encoder in the second stage.
-For example, the AudioCaps R@1 performance improved from M2D2-’s 27.34 to M2D2’s 37.24.
-The results demonstrate the importance of training the text encoder to achieve fine-grained alignment of the paired features.
-The results in Table VIII indicate that the combination of datasets significantly impacts performance.
-The “Final CLAP training dataset” column in the table lists the datasets used in the model’s last training just before the evaluation.
-LAION-CLAP, CLAP2023, and M2D2, which used a large-scale dataset (denoted as ”LS” in the table) just before the evaluation, perform poorly compared to those without one.
+Table VIII shows that M2D-CLAP achieves top performance, and the ablation variants improve the performance as stages go by.
+The first-stage model, M2D-CLAP, exhibits a large gap from the previous SOTA results, such as HTSAT-BERT-FT.
+The second-stage pre-training significantly improves the performance, e.g., from 27.34 of the first stage to 37.24 of the second stage for AudioCaps Text-to-Audio R@1.
+This indicates that pre-training the text encoder is crucial for a task that requires finer alignment between audio and text, unlike the zero-shot classification, where the first-stage model achieved comparable results.
+The additional stages 1.1 (AudioSet fine-tuning) and 2.1 (WavCaps fine-tuning) bring the final performance to the top level, and M2D-CLAP even outperforms top performers HTSAT-BERT-FT, Cacophony, and FLAP in AudioCaps.
+M2D-CLAP outperforms both M2D-CLAP and M2D-CLAP+WC.
+Notably, the latter is pre-trained on the same dataset as M2D-CLAP, ensuring that the only differences lie in the text encoder and audio projector.
+M2D-CLAP improves performance on most metrics, e.g., AudioCaps Audio-to-Text R@1 performance of 36.05 compared to M2D-CLAP+WC’s 28.63.
+The results confirm the effectiveness of using an LLM-based text (sentence) encoder and a transformer-based audio projector.
+Table VIII indicates that the combination of datasets significantly impacts performance.
+The “Final CLAP training dataset” column lists the datasets used in the model’s last training just before the evaluation.
+LAION-CLAP, CLAP2023, and M2D-CLAP, which used a large-scale dataset (denoted as ”LS” in the table), perform poorly compared to those without one.
 Their AudioCaps Text-to-Audio R@1 performance is about 35 to 37, which is inferior to the 39.7 of HTSAT-BERT-PT without using large-scale datasets, for example.
@@
 As an exception, Cacophony, which developed its data curation strategy, performs well with large-scale datasets.
-M2D2+ and M2D2 achieve top performance when additionally trained under conditions advantageous to the benchmark as in the previous studies[4][6][7].
-M2D2+ outperforms or becomes comparable to HTSAT-BERT-PT after pre-training on the same data (WavCaps, AudioCaps, and Clotho) as HTSAT-BERT-PT in additional stage 2.1.
-M2D2 even outperforms top performers HTSAT-BERT-FT, Cacophony, and FLAP in AudioCaps after fine-tuning its audio encoder to AudioSet in the additional stage 1.1.
-In summary, the experiments validated the effectiveness of M2D2.
-M2D2 performs comparably to previous studies, and it further improves performance to the top level through additional stages for a fair comparison to previous studies.
+In summary, the experiments validated the effectiveness of M2D-CLAP and demonstrated that the staged training process enables the model to achieve fine-grained alignment between paired features.
 | Audio enc. AudioSet fine-tuning | AudioCaps (subset of AudioSet) | Clotho (out of AudioSet) | |||||||||
@@
 | CLAP2023[6] | ✓ | 22.4 0.5 | 65.2 2.1 | 16.3 0.3 | 40.7 1.2 | 60.5 0.8 | 17.5 0.1 | 42.5 0.5 | 12.3 0.1 | 27.4 0.3 | 49.2 0.2 |
-| (Proposed models) | |||||||||||
-| M2D2- | - | 23.3 0.3 | 68.3 1.5 | 16.9 0.2 | 42.6 0.8 | 61.9 0.5 | 17.4 0.1 | 41.2 0.3 | 12.2 0.1 | 26.7 0.2 | 48.9 0.2 |
-| M2D2 | - | 23.7 0.3 | 69.6 1.3 | 17.2 0.3 | 43.4 0.7 | 62.5 0.4 | 17.8 0.1 | 43.5 0.4 | 12.4 0.1 | 28.0 0.2 | 49.9 0.2 |
-| M2D2+ | - | 23.6 0.3 | 69.4 1.4 | 17.1 0.3 | 43.2 0.8 | 62.3 0.5 | 17.8 0.1 | 43.2 0.4 | 12.5 0.1 | 27.8 0.3 | 49.7 0.2 |
-| M2D2 | ✓ | 24.3 0.4 | 72.4 1.8 | 17.6 0.3 | 45.0 1.0 | 63.9 0.4 | 17.7 0.1 | 43.1 0.7 | 12.5 0.2 | 27.8 0.4 | 49.6 0.4 |
+| (Proposed model) | |||||||||||
+| M2D-CLAP | ✓ | 24.3 0.4 | 72.4 1.8 | 17.6 0.3 | 45.0 1.0 | 63.9 0.4 | 17.7 0.1 | 43.1 0.7 | 12.5 0.2 | 27.8 0.4 | 49.6 0.4 |
+| M2D-CLAP | - | 23.3 0.3 | 68.3 1.5 | 16.9 0.2 | 42.6 0.8 | 61.9 0.5 | 17.4 0.1 | 41.2 0.3 | 12.2 0.1 | 26.7 0.2 | 48.9 0.2 |
+| M2D-CLAP | - | 23.7 0.3 | 69.6 1.3 | 17.2 0.3 | 43.4 0.7 | 62.5 0.4 | 17.8 0.1 | 43.5 0.4 | 12.4 0.1 | 28.0 0.2 | 49.9 0.2 |
+| M2D-CLAP | - | 23.6 0.3 | 69.4 1.4 | 17.1 0.3 | 43.2 0.8 | 62.3 0.5 | 17.8 0.1 | 43.2 0.4 | 12.5 0.1 | 27.8 0.3 | 49.7 0.2 |
 | (Reference results from non-EnCLAP models) | |||||||||||
@@
 We used the same settings for other hyperparameters as in EnCLAP, such as training epochs of 15.
-We tested the top-five checkpoints of SPIDEr validation performance on the test set and averaged the results to obtain the final result.
+We tested the top five checkpoints of SPIDEr validation performance on the test set and averaged the results to obtain the final result.
 The evaluation metrics are METEOR[85], CIDEr[86], SPICE[87], SPIDEr[88], and FENSE[89].
 Results: Audio captioning
-Experimental results in Table IX show that M2D2 is comparable to LAION-CLAP on AudioCaps, while M2D2, M2D2+, and M2D2 outperform LAION-CLAP on Clotho.
+Table IX shows that M2D-CLAP archives top performance, which is comparable to LAION-CLAP on AudioCaps and better than LAION-CLAP on Clotho.
 We also evaluated CLAP2023; however, it underperformed LAION-CLAP.
-Note that the experiments used the same EnCLAP codebase; the only difference is the CLAP audio feature used by the EnCLAP.
-The AudioCaps results demonstrate that fine-tuning the audio encoder to AudioSet is vital for good performance; M2D2 performs comparably to LAION-CLAP while M2D2+ does not.
-On the other hand, the Clotho results suggest that fine-tuning to AudioSet does not offer general-purpose performance; M2D2 and M2D2+ perform comparably to M2D2, and they all outperform LAION-CLAP.
-Notably, M2D2 performs slightly better than M2D2+ and M2D2 on Clotho, demonstrating that it learns generalizable CLAP features without the need for benchmark techniques.
-In summary, the experimental results confirm that the M2D2 models achieve CLAP features that also perform well in audio captioning.
-IV-D M2D2 Analysis
+Note that the experiments used the same EnCLAP codebase; the only difference is the CLAP audio feature used by EnCLAP.
+The AudioCaps results of intermediate stages, M2D-CLAP, underperform those of LAION-CLAP, indicating that fine-tuning the audio encoder to AudioSet is vital for good performance with AudioCaps.
+On the other hand, the Clotho results suggest that fine-tuning to AudioSet does not offer general-purpose performance; all M2D-CLAP variants outperform LAION-CLAP.
+Notably, M2D-CLAP performs slightly better than M2D-CLAP on Clotho, demonstrating that it learns generalizable CLAP features without the need for fine-tuning to AudioSet and WavCaps.
+In summary, the experimental results confirm that M2D-CLAP produces CLAP features that also perform well in audio captioning.
+IV-D M2D-CLAP Analysis
 IV-D1 Taming the CLAP objective in the first stage
@@
 4 shows the average results of linear evaluation on general tasks across training epochs for (a) all tasks, (b) ESC-50/US8K tasks for which the CLAP objective is effective, and (c) all other tasks when the weight is varied.
-When the weight is at 0.1, (b) ESC-50/US8K performance matures quickly and degrades as training progresses, while (c) performance for other tasks does not improve.
-When the weight is at 0.001, (b) ESC-50/US8K performance improves slowly, suggesting the weight is too low.
-In contrast, with the weight at 0.01, performance for all tasks matures as training progresses.
-The ESC-50/US8K performance’s maturing quickly with high weight is similar to the observation that the CLAP training matures early in the previous study[6].
-The observations of all cases suggest that combining SSL and the CLAP objectives while suppressing CLAP is essential for successful pre-training.
+With a weight of 0.1, (b) ESC-50/US8K performance matures quickly but degrades later, while (c) other tasks show no improvement.
+At 0.001, (b) performance improves slowly, indicating the weight is too low.
+In contrast, 0.01 allows (a) all tasks to mature steadily during training.
+For reference, at weight 0.0 (i.e., M2D), (b) is low due to lack of caption supervision, (c) remains high, and (a) is slightly lower overall than M2D-CLAP.
+The rapid maturation of ESC-50/US8K performance with higher CLAP weights mirrors the early maturation of CLAP training reported in prior work[6].
+These results suggest that suppressing the effect of CLAP while combining it with SSL is key for effective pre-training.
 Audio projector ablations
-The audio projector, which aggregates the audio features of each token into a single feature, impacts the effectiveness of training through CLAP loss calculation.
+The audio projector, which aggregates token-level audio features into a single feature for CLAP loss computation, impacts training effectiveness.
 Fig.
-5 shows the average results of linear evaluation across training epochs for the ablations of audio projectors: M2D2’s one-block transformer encoder, the two-block encoder (one additional block), and one-hidden-layer MLP used in M2D-CLAP (M2D20-).
-When using MLP as a projector, we averaged all the audio encoder output features and fed the averaged feature to it.
-The results in Fig.
-5(a) show that MLP exhibits the best results up to 200 epochs but stalls afterward.
-In the final 300 epochs, one block transformer of M2D2 outperforms the others and obtains the best results.
-The MLP results in Fig.
-5(b) show a trend similar to the ESC-50/US8K results when the CLAP loss weight is high () in Fig.
-4(b), suggesting that the use of a simple MLP augments the contribution of the CLAP objective during training and leads to early maturation, as reported in the previous study[6].
-The two-block transformer underperformed the one-block transformer throughout the training process, indicating that using a projector with adequate network capacity is also vital for successful pre-training.
+5 compares average linear evaluation results across epochs for different audio projectors: the one-block transformer in M2D-CLAP, a two-block transformer (with one additional block), and a one-hidden-layer MLP used in the previous M2D-CLAP[11].
+For the MLP, we input the average of all audio encoder outputs.
+In Fig.
+5(a), the MLP performs best up to 200 epochs but stalls afterward.
+In the final 300 epochs, M2D-CLAP’s one-block transformer surpasses the others.
+The MLP’s trend (Fig.
+5(b)) resembles ESC-50/US8K results under a high CLAP loss weight () in Fig.
+4(b), suggesting that a simple MLP increases CLAP’s influence during training and leads to early maturation, in line with previous findings[6].
+The two-block transformer consistently underperforms the one-block version, indicating that moderate network capacity is important for effective pre-training.
 IV-D2 Text encoder ablations
-We assessed the performance impact on the type of text encoder used in the first stage.
-We compared the LLM-based model NV-Embed-v2[21][42] (7.85B parameters) used in M2D2 against the GTE-Base[60] (109M parameters) used in M2D-CLAP (M2D20-) and GTE-Qwen2-7B-instruct999https://huggingface.co/Alibaba-NLP/gte-Qwen2-7B-instruct (7.61B parameters)[60], which ranked higher than NV-Embed-v2 on the MTLB leaderboard.
-The experiments replaced the text encoder in M2D2.
+We assessed the performance impact of the type of text encoder used in the first stage.
+We compared the LLM-based model NV-Embed-v2[21][42] (7.85B parameters) used in M2D-CLAP against the GTE-Base[64] (109M parameters) used in the previous M2D-CLAP (M2D-CLAP) and GTE-Qwen2-7B-instruct999https://huggingface.co/Alibaba-NLP/gte-Qwen2-7B-instruct (7.61B parameters)[64], which ranked higher than NV-Embed-v2 on the MTLB leaderboard.
+The experiments replaced the text encoder in M2D-CLAP.
 We also assessed the performance of the second-stage text encoder when using GTE-Base instead of BERT.
 The linear evaluation results on the general audio tasks in Table X and the audio-text retrieval of the CLAP features in Table XI show that using NV-Embed-v2’s LLM-based sentence features achieves the best performance with both features.
-With a non-LLM-based GTE-Base, M2D2 learns competitive audio features while underperforming second-stage CLAP features compared to LLM-based encoders.
+With a non-LLM-based GTE-Base, M2D-CLAP learns competitive audio features while underperforming second-stage CLAP features compared to LLM-based encoders.
 The GTE-Qwen2-7B-instruct’s results are comparable to (while slightly worse than) NV-Embed-v2’s, confirming the effectiveness of employing LLM-based text encoders.
@@
 | Bold results are better than the NV-Embed-v2 (NVEv2BERT) results above. | ||||||||
-Notably, the performance boost in the second-stage CLAP features indicates that the semantic knowledge accumulated in the audio features effectively guides M2D2 to learn useful CLAP features.
+Notably, the performance boost in the second-stage CLAP features indicates that the semantic knowledge accumulated in the audio features effectively contributes to learn useful CLAP features.
 The second-stage results of LLM-based NV-Embed-v2 and GTE-Qwen2-7B-instruct are 37.24 and 36.78 for R@1 of AudioCaps T-to-A, while the non-LLM-based GTE-Base’s result is 34.04, for example.
@@
 IV-D3 Training design ablations
-Training audio encoder in the second stage
-While M2D2 freezes the audio encoder in the second stage, an alternative would be to pre-train the audio and text encoder together in this stage, as in [10].
-We performed an ablation experiment in which we pre-trained the audio encoder in the second stage.
-In this case, computational resource constraints decreased the batch size from 2048 to 1024.
-We also compared the results with M2D2, in which we fine-tuned the audio encoder to AudioSet after the first stage, i.e., the additional stage 1.1 described in Section III-B4.
-The results for the audio features shown in Table XII indicate that M2D2, which trains the audio encoder in the second stage, improves the performance of the audio features in the ESC-50 (ESC) and US8K tasks while degrading the performance in other tasks.
-This trend is the same as M2D2.
-On the other hand, Table XIII shows that the audio-text retrieval performance of M2D2 is comparable to M2D2, indicating that training the audio encoder in the second stage does not improve the CLAP feature, which reveals the difference from M2D2.
-In addition, the decreased batch size may have affected the results.
-These experiments suggest that training audio encoders in the second stage, as in the previous study Cacophony[10], might not be a good pre-training strategy for learning representations that perform well for both audio and CLAP features, the goal of M2D2.
+We conducted the following ablations focusing on differences from conventional CLAP in the M2D-CLAP training process:
+1.
+Training audio encoder in the second stage: While M2D-CLAP freezes the audio encoder in the second stage, an alternative is to pre-train both audio and text encoders, as in [10].
+This ablation adopts that approach, while reducing the batch size to 1024 from 2048 due to limited computational resources.
+2.
+Using text projector only: This ablation removed the audio projector and added a text projector on top of the text encoder, aiming to learn an audio feature directly in the semantic space that serves both conventional and CLAP audio features.
+We used a one-hidden-layer MLP as the text projector.
+3.
+Adding text projector: This ablation added a text projector on top of the text encoder, making M2D-CLAP follow a typical CLAP configuration with both audio and text projectors.
+Tables XII and XIII present the evaluation results for audio and CLAP features, respectively, with key observations summarized below:
+(i) Training audio encoder in the second stage: For audio features, the ablation (Stage 2/TA) improves performance on ESC-50 (ESC) and US8K, but clearly degrades it on speech tasks.
+This trend, similar to that of the final M2D-CLAP, suggests an influence of caption supervision.
+For CLAP features, Stage 2/TA lowers A-to-T performance in both retrieval tasks.
+Unlike previous studies, factors such as pre-training the audio encoder under M2D and CLAP objectives, omitting the text projector, or using a non-MLP audio projector may have slightly hindered fine-grained audio-to-text alignment.
+Additionally, the reduced batch size may have negatively impacted results.
+(ii) Using text projector only: Unlike other ablations, the ablations (Stages 1, 2/TP only) degrade performance across all results for both features.
+Thus, these results highlight the challenge of bridging conventional and CLAP audio features.
+(iii) Adding text projector: The ablation (Stages 1/+TP) causes minimal changes in audio feature performance, while Stages 2/+TP clearly degrades CLAP feature performance.
+Mapping text features through a projector appears to hinder overall audio-text alignment in CLAP features.
+In summary, unlike previous work such as Cacophony[10], continuing to train the audio encoder in Stage 2 is ineffective, and incorporating a text projector yields no clear advantage for M2D-CLAP.
 | Model | ESC | US8K | SPC | VC1 | VF | CRD | GTZ | NS | SUR | Avg. |
-| M2D2, M2D2- | 95.2 | 89.5 | 95.6 | 68.6 | 98.1 | 73.4 | 86.6 | 76.5 | 42.2 | 80.6 |
-| M2D2, M2D2 | 97.9 | 89.7 | 94.8 | 56.3 | 97.1 | 69.7 | 86.3 | 76.7 | 41.1 | 78.8 |
-| (Section IV-D3: Jointly training text/audio encoders in the second stage) | ||||||||||
-| M2D2 | 97.0 | 90.1 | 93.3 | 60.4 | 96.7 | 71.1 | 86.3 | 76.1 | 42.0 | 79.0 |
-| (Section IV-D3: Text projector only) | ||||||||||
-| M2D2-(TP only) | 94.9 | 88.0 | 94.0 | 60.5 | 97.5 | 71.6 | 83.7 | 76.0 | 41.9 | 78.7 |
-| (Section IV-D3: Both audio and text projector, as in previous CLAP models) | ||||||||||
-| M2D2-(+TP) | 96.1 | 88.9 | 95.6 | 66.2 | 97.8 | 72.7 | 84.7 | 77.6 | 42.3 | 80.2 |
+| Stage 1 | 95.2 | 89.5 | 95.6 | 68.6 | 98.1 | 73.4 | 86.6 | 76.5 | 42.2 | 80.6 |
+| (i) Training audio encoder in the second stage. | ||||||||||
+| Stage 2/TA | 97.0 | 90.1 | 93.3 | 60.4 | 96.7 | 71.1 | 86.3 | 76.1 | 42.0 | 79.0 |
+| (ii) Using text projector only. | ||||||||||
+| Stage 1/TP only | 94.9 | 88.0 | 94.0 | 60.5 | 97.5 | 71.6 | 83.7 | 76.0 | 41.9 | 78.7 |
+| (iii) Adding text projector, as in previous CLAP models. | ||||||||||
+| Stage1/+TP | 96.1 | 88.9 | 95.6 | 66.2 | 97.8 | 72.7 | 84.7 | 77.6 | 42.3 | 80.2 |
+| M2D-CLAP | 97.9 | 89.7 | 94.8 | 56.3 | 97.1 | 69.7 | 86.3 | 76.7 | 41.1 | 78.8 |
 | AudioCaps | Clotho | |||||||
@@
 | R@1 | mAP @10 | R@1 | mAP @10 | R@1 | mAP @10 | R@1 | mAP @10 | |
-| M2D2 | 37.24 | 52.00 | 52.56 | 38.74 | 18.09 | 28.51 | 24.88 | 15.53 |
-| M2D2 | 39.06 | 54.29 | 54.02 | 39.95 | 18.30 | 29.15 | 22.78 | 15.62 |
-| (Section IV-D3: Jointly training text/audio encoders in the second stage) | ||||||||
-| M2D2 | 37.81 | 52.29 | 50.78 | 37.15 | 17.76 | 28.83 | 21.44 | 14.11 |
-| (Section IV-D3: Text projector only) | ||||||||
-| M2D2 (TP only) | 34.04 | 48.73 | 45.98 | 34.42 | 17.07 | 27.44 | 22.11 | 14.56 |
-| (Section IV-D3: Both audio and text projector, as in previous CLAP models) | ||||||||
-| M2D2 (+TP) | 36.07 | 50.52 | 48.69 | 36.93 | 17.53 | 28.09 | 22.58 | 14.72 |
-Using text projector only
-In this ablation experiment, we removed the audio projector and attached a text projector on top of the text encoder to test the case where an audio feature is learned in the semantic space, and we mapped text features to the space, i.e., an audio feature was learned that works as both conventional and CLAP audio features.
-We used a one-hidden layer MLP as the text projector.
-Adding text projector
-This ablation experiment added a text projector to the text encoder, making M2D2 follow a typical CLAP configuration with both audio and text projectors.
-The results in Tables XII and XIII show that M2D2 (+TP), text projector-added model, degrades performance from M2D2 as in the text projector only ablation.
-Considering these results, we do not employ a text projector in M2D2.
+| Stage 2 | 37.24 | 52.00 | 52.56 | 38.74 | 18.09 | 28.51 | 24.88 | 15.53 |
+| (i) Training audio encoder in the second stage. | ||||||||
+| Stage 2/TA | 37.81 | 52.29 | 50.78 | 37.15 | 17.76 | 28.83 | 21.44 | 14.11 |
+| (ii) Using text projector only. | ||||||||
+| Stage 2/TP only | 34.04 | 48.73 | 45.98 | 34.42 | 17.07 | 27.44 | 22.11 | 14.56 |
+| (iii) Adding text projector, as in previous CLAP models. | ||||||||
+| Stage 2/+TP | 36.07 | 50.52 | 48.69 | 36.93 | 17.53 | 28.09 | 22.58 | 14.72 |
+| M2D-CLAP | 39.06 | 54.29 | 54.02 | 39.95 | 18.30 | 29.15 | 22.78 | 15.62 |
 IV-D4 Visualizations
@@
 We visualize the pre-trained CLAP audio features and show their impact on two downstream tasks with an inherently different nature.
-We encoded ESC-50 and VoxCeleb1 sounds using M2D2, M2D2 (M2D2 fine-tuned to AudioSet), and CLAP and visualized their features using t-SNE as shown in Fig.
+We encoded ESC-50 and VoxCeleb1 sounds using M2D-CLAP and CLAP and visualized their features using t-SNE as shown in Fig.
 6.
 We use all 50 class samples from ESC-50 and samples of random ten speakers from VoxCeleb1.
-The ESC-50 results show that for all audio features, the samples are clustered by class, indicating that the features are useful for the 50-class classification task.
-On the other hand, the VoxCeleb1 results show that the CLAP clusters have unclear boundaries compared to M2D2 and M2D2.
-These observations support the VC1 performance difference in the linear evaluation shown in Table II in Section IV-B1.
+We also computed the silhouette score to quantitatively assess the quality of clusters formed by each model’s features.
+The ESC-50 results show that both model features form clear class-wise clusters with high silhouette scores, consistent with their strong performance in the linear evaluation results in Table II in Section IV-B1.
+On the other hand, the VoxCeleb1 results show that CLAP forms slightly less distinct clusters (silhouette score of 0.072) compared to M2D-CLAP (silhouette score of 0.109).
+These observations support the VC1 performance difference in Table II.
 The CLAP features are advantageous for ESC-50 since it is an environmental sound classification task, and the pre-training captions contain much information indicating differences in environmental sounds.
@@
 The conventional CLAP model would learn to represent speaker attributes such as adult or child described in captions, which may explain the tendency not to cluster by individuals.
-In contrast, M2D2 features form more distinct clusters, indicating that SSL helps learn more information about sound itself, which leads to clearer boundaries for individuals.
-These observations demonstrate that jointly learning SSL and CLAP allows M2D2 to include diverse information in CLAP audio features.
+In contrast, M2D-CLAP features form more distinct clusters, indicating that SSL helps learn more descriptive information about sound itself, which leads to clearer clusters for individuals.
+These observations demonstrate that jointly learning SSL and CLAP allows M2D-CLAP to encode both discriminative and descriptive information in CLAP audio features.
 Audio projector attention visualizations
-Figure 7 shows the attention in the audio projector for the audio samples from AudioCaps and Clotho.
+Figure 7 shows the attention in the audio projector in M2D-CLAP for the audio samples from AudioCaps and Clotho.
 The attention map directly shows which patch features the final CLAP audio feature attends to because the audio projector is a single-layer transformer encoder with a single head.
-We used the M2D2 model and picked the sounds with typical attention maps.
+We picked the sounds with typical attention maps.
 The first example from AudioCaps shows that the projector attends to the louder part of the revving sound and, notably, to the sound characterizing the powering down.
@@
 Overall, the observations exemplify that the projector summarizes CLAP audio features based on the patch features that characterize the audio clips.
+IV-D5 Potential and limitations
+The application of top-performing audio features to medical (respiratory and heart) sounds in [90] demonstrates their practical utility, suggesting broader potential for real-world problem solving.
+CLAP features have also been applied to diverse tasks, such as music generation[91] and foley sound synthesis[92].
+Thus, M2D-CLAP has the potential to contribute to a wide range of applications by providing both general audio and CLAP features.
+On the other hand, it still has limitations in extending its applicability, particularly regarding time resolution and frequency range.
+It processes sounds with a temporal resolution of 160 ms and a sampling rate of 16 kHz, and adapting it to tasks with different requirements would necessitate retraining.
+While effectiveness has been shown even with limited frequency content, such as 4 kHz audio in [90], extending the approach to more diverse music tasks requiring a wider frequency range, as well as to the bioacoustic domain requiring focus on specific frequency ranges depending on the task, remains a future research direction.
+The CLAP features of M2D-CLAP also have a limitation in the data coverage for general-purpose usage.
+Achieving general-purpose performance, as in cutting-edge LLM, may require a similar scale of training data.
+To this end, further advancement of approaches such as automated generation of audio captions, as in Auto-ACD[46], is necessary to address the challenge of data scalability.
 V Conclusion
 This study pursued a general-purpose audio-language representation that provides both general-purpose audio and CLAP features to serve diverse audio applications.
-For this purpose, we proposed M2D2, a new method combining an SSL M2D and CLAP in a two-stage pre-training and utilizing a recent LLM-based sentence encoder.
-In the first stage, M2D2 learns a generalizable audio feature using the rich semantic features of the LLM-based sentence encoder, and in the second stage, it learns a CLAP feature using the learned audio feature.
-While M2D learns highly generalized audio features, the powerful LLM-based embeddings help learn a better CLAP feature through the audio feature.
-Experiments validated the effectiveness of M2D2 as a general-purpose audio-language representation, highlighted by the SOTA performance of audio features for AudioSet fine-tuning and music tasks and top performance of CLAP features in zero-shot classification, audio-text retrieval, and audio captioning.
-We conducted ablation studies to analyze M2D2 pre-training and found that two-stage training strategy design and managing the effect of the CLAP objective are vital for learning effective features.
-We have made our models available online and hope they help the progress of diverse audio studies101010To reviewers, we will update our repository after the notification..
+For this purpose, we proposed M2D-CLAP, a novel method combining an SSL M2D and CLAP in a multi-stage pre-training and utilizing a recent LLM-based sentence encoder.
+In the first stage, M2D-CLAP learns a generalizable audio feature using the rich semantic features of the LLM-based sentence embedding model, and in the second stage, it learns a CLAP feature using the learned audio feature.
+It further refines features by fine-tuning.
+Experiments validated the effectiveness of M2D-CLAP as a general-purpose audio-language representation.
+Its audio features performed well, particularly in AudioSet fine-tuning and music tasks, and its CLAP features performed well in zero-shot classification, audio-text retrieval, and audio captioning.
+These results demonstrate the novelty of M2D-CLAP in providing both high-performing general audio and CLAP features.
+We conducted ablation studies to analyze M2D-CLAP and found that a multi-stage training strategy design and managing the effect of the CLAP objective are vital for learning effective features.
+We release our models to advance diverse audio studies.